### v1

# FC-Attack: Jailbreaking Large Vision-Language Models via Auto-Generated Flowcharts

Abstract
Large Vision-Language Models (LVLMs) have become powerful and widely adopted in some practical applications. However, recent research has revealed their vulnerability to multimodal jailbreak attacks, whereby the model can be induced to generate harmful content, leading to safety risks. Although most LVLMs have undergone safety alignment, recent research shows that the visual modality is still vulnerable to jailbreak attacks.
In our work, we discover that by using flowcharts with partially harmful information, LVLMs can be induced to provide additional harmful details. Based on this, we propose a jailbreak attack method based on auto-generated flowcharts, $\mathsf{FC\text{-}Attack}$. Specifically, $\mathsf{FC\text{-}Attack}$ first fine-tunes a pre-trained LLM to create a step-description generator based on benign datasets. The generator is then used to produce step descriptions corresponding to a harmful query, which are transformed into flowcharts in $3$ different shapes (vertical, horizontal, and S-shaped) as visual prompts. These flowcharts are then combined with a benign textual prompt to execute a jailbreak attack on LVLMs. Our evaluations using the Advbench dataset show that $\mathsf{FC\text{-}Attack}$ achieves over $90\%$ attack success rates on Gemini-1.5, Llaval-Next, Qwen2-VL, and InternVL-2.5 models, outperforming existing LVLM jailbreak methods. Additionally, we investigate factors affecting the attack performance, including the number of steps and the font styles in the flowcharts. Our evaluation shows that $\mathsf{FC\text{-}Attack}$ can improve the jailbreak performance from $4\%$ to $28\%$ in Claude-3.5 by changing the font style.
To mitigate the attack, we explore several defenses and find that AdaShield can largely reduce the jailbreak performance but with the cost of utility drop.
Disclaimer: This paper contains examples of harmful language. Reader discretion is recommended.
Figure 1: Comparison of jailbreak effectiveness in various LVLMs using three prompt types: (1) harmful text without images; (2) harmful step descriptions without images; and (3) harmful step descriptions in images.
Introduction
With the advancement of Large Language Models (LLMs), Large Vision-Language Models (LVLMs) that integrate vision and text, such as GPT-4o [14], have demonstrated emergent abilities and achieved impressive performance on downstream tasks related to visual understanding [20, 16].
Despite being powerful, recent studies [11, 27] have revealed that LVLMs are vulnerable to jailbreak attacks whereby the adversary uses malicious methods to bypass safeguards in LVLMs and gain harmful knowledge. Such vulnerabilities pose remarkable safety risks to the Internet and the physical world. For instance, in January 2025, the world witnessed the first case where ChatGPT was used to conduct an explosion. [^1] https://www.thetimes.com/world/us-world/article/vegas-cybertruck-bomber-used-chatgpt-to-plan-explosion-wk7h2jtkq. To better safeguard LVLMs and proactively address their vulnerabilities, model researchers make many efforts in this regard, such as Zhao et al. [42] providing a quantitative understanding regarding the adversarial vulnerability of LVLMs. Previous studies often create adversarial datasets tailored to specific models, which tend to perform poorly on other models.
Currently, jailbreak attacks against LVLMs can be broadly categorized into two main types: optimization-based attacks [5, 19] and prompt-based attacks [11, 33]. Optimization-based attacks involve identifying an image that semantically corresponds to a harmful textual prompt. Adversarial perturbations are then added to this image using gradient-based optimization techniques to achieve the jailbreak. These methods usually require white-box access and are relatively slow due to the optimization process. In contrast, prompt-based jailbreak methods require only black-box access, which can only observe the model’s outputs without direct access to its internal parameters. These attacks exploit the vulnerabilities in LVLMs’ alignment mechanisms, which focus mainly on text-based safety alignment. By injecting malicious information into the visual prompt, these methods guide LVLMs in generating harmful content. Generally, optimization-based methods are more effective in white-box scenarios but show limited transferability to black-box models. Prompt-based attacks are relatively straightforward. For example, FigStep merely embeds harmful queries into images without employing additional attack vectors, facilitating simpler defense against these attacks.
To better improve the attack transferability and its effectiveness, we propose a novel prompt-based jailbreak attack, namely $\mathsf{FC\text{-}Attack}$. Concretely, $\mathsf{FC\text{-}Attack}$ converts harmful queries into harmful flowcharts as visual prompts, allowing users to input benign textual prompts to bypass the model’s safeguards. Specifically, $\mathsf{FC\text{-}Attack}$ consists of two stages: (1) Step-Description Generator Building: In this stage, the step description dataset is synthesized using GPT-4o, and Low-Rank Adaptation (LoRA) is applied to fine-tune a pre-trained LLM to obtain a step-description generator. (2) Jailbreak Deployment: This stage uses the generator mentioned above to produce execution steps corresponding to the harmful query and generates three types of harmful flowcharts (vertical, horizontal, and S-shaped) as visual prompts. Together with the benign textual prompt, the visual prompt will be fed into LVLM to achieve the jailbreak. Note that the harmful flowcharts are generated automatically without hand-crafted effort, which enhances their efficiency.
Our evaluation on the Advbench dataset shows that $\mathsf{FC\text{-}Attack}$ outperforms previous attacks and achieves an attack success rate (ASR) of over $90\%$ on multiple open-source models, including Llava-Next, Qwen2-VL, and InternVL-2.5, and reaches $94\%$ on the production model Gemini-1.5. Although the ASR is lower on GPT-4o mini, GPT-4o, and Claude-3.5, we how later that it can be improved in certain ways. To further investigate the impact of different elements in flowcharts on the jailbreak effectiveness of LVLMs, we conduct several ablation experiments, including different types of user queries (as shown in Figure 1), numbers of descriptions, and font styles in flowcharts. These experiments show that LVLMs exhibit higher safety in the text modality but weaker in the visual modality. Moreover, we find that even flowcharts with a one-step harmful description can achieve high ASR, as evidenced by the Gemini-1.5 model, where the ASR reaches $86\%$. Furthermore, font styles in flowcharts also contribute to the ASR increase. For instance, when the font style is changed from “Times New Roman” to “Pacifico”, the ASR increases from $4\%$ to $28\%$ on the model with the lowest ASR (Claude-3.5) under the original style. To mitigate the attack, we consider several popular defense approaches, including Llama-Guard-3-11B-Vision [24], JailGuard [39], AdaShield-S [32], and AdaShield-A [32]. Among them, AdaShield-A demonstrates the best defense performance by reducing the average ASR from $58.6\%$ to $1.7\%$. However, it also reduces LVLM’s utility on benign datasets, which calls for more effective defenses.
Overall, our contributions are summarized as follows:
  • We propose a novel method, $\mathsf{FC\text{-}Attack}$ that jailbreaks LVLMs using auto-generated harmful flowcharts. Experiments conducted on Advbench show that $\mathsf{FC\text{-}Attack}$ achieves a higher ASR across multiple models compared to other LVLM jailbreak attacks.
  • Our ablation study investigates the impact of different types of user queries, the number of steps, and the font style in flowcharts. We find that the font style could serve as a key factor to further improve the ASR, especially for safer LVLMs, revealing a novel attack channel in LVLMs.
  • We explore multiple defense strategies and find that AdaShield-A effectively reduces the ASR of $\mathsf{FC\text{-}Attack}$, but with the cost of reducing model utility.
Related Work
Large Vision-Language Model
In recent years, with the increase in model parameters and training data, LLMs have demonstrated powerful language generation and understanding capabilities [41, 7], which have driven the emergence of LVLMs [38] (also known as Multimodal Large Language Models, MLLMs). LVLMs combine visual understanding with language comprehension, showing promising capabilities in visual downstream tasks, including Visual Question Answering (VQA) [3, 17, 28], image captioning [13, 18], and visual commonsense reasoning [37, 30].
In this paper, we consider both popular open-source and widely used production LVLMs. These LVLMs are the most widely used, and all of them have been aligned to ensure safety. Detailed information are introduced in Appendix A
Figure 2: Overview of the $\mathsf{FC\text{-}Attack}$ framework, including two stages: Step-Description Generator Building and Jailbreak Deployment.
Jailbreak Attacks on LVLMs
Similar to LLMs, despite undergoing safety alignment, LVLMs remain susceptible to jailbreak attacks. Current attacks can be categorized into two types: optimization-based and prompt-based attacks (also known as non-optimization-based attacks).
Most existing optimization-based attacks rely on backpropagating the gradient of the target to generate harmful outputs. These methods typically require white-box access to the model, where they obtain the output logits of LVLMs and then compute the loss with the target response to create adversarial perturbations into the visual prompts or textual prompts  [4, 29, 26] (e.g., the target can be “Sure! I’m ready to answer your question.”). Carlini et al. [6] are the first to propose optimizing input images by using fixed toxic outputs as targets, thereby forcing the model to produce harmful outputs. Building on this, Bailey et al. [5] introduce the Behaviour Matching Algorithm, which trains adversarial images to make LVLMs output behavior that matches a target in specific contextual inputs. This process requires the model’s output logits to align closely with those of the target behavior. Additionally, they propose Prompt Matching, where images are used to induce the model to respond to specific prompts. Li et al. [19] take this further by replacing harmful keywords in the original textual inputs with objects or actions in the image, allowing harmful information to be conveyed through images to achieve jailbreaking. Unlike previous work, these images are generated using diffusion models and are iteratively optimized with models like GPT-4. This approach enhances the harmfulness of the images, enabling more effective attacks.
Unlike optimization-based attacks, prompt-based attacks only need black-box access to successfully attack the model without introducing adversarial perturbations into images. Gong et al. [11] discovers that introducing visual modules may cause the original security mechanisms of LLMs to fail in covering newly added visual content, resulting in potential security vulnerabilities. To address this, they propose the FigStep attack, which converts harmful textual instructions into text embedded in images and uses a neutral textual prompt to guide the model into generating harmful content. This method can effectively attack LVLMs without requiring any training. Wang et al. [33] identifies a phenomenon called Shuffle Inconsistency, which highlights the tension between “understanding capabilities” and “safety mechanisms” of LLMs. Specifically, even if harmful instructions in text or images are rearranged, LVLMs can still correctly interpret their meaning. However, the safety mechanisms of LVLMs are often more easily bypassed by shuffled harmful inputs than by unshuffled ones, leading to dangerous outputs. Compared to optimization-based attacks, prompt-based attacks are simple to achieve high success rates against closed-source models. Our proposed $\mathsf{FC\text{-}Attack}$ also belongs to this category, requiring only black-box access to successfully jailbreak the model.
Threat Model
Adversary’s Goal
The adversary’s goal is to exploit attack methods to bypass the protective mechanisms of LVLMs and access content prohibited by safety policies, e.g., OpenAI’s usage policy. [^2] https://openai.com/policies/usage-policies This goal takes real-world scenarios into account, where adversaries manipulate the capabilities of powerful LVLMs to easily acquire harmful knowledge and thereby commit criminal acts with minimal learning effort. These objectives pose severe societal impacts and risks, turning model providers into indirect accomplices, and thereby damaging their reputations.
Adversary’s Capabilities
In this paper, we consider a black-box scenario where the adversary cannot directly access the model’s structure, parameters, or output logits, but can only obtain the model’s final output (texts). In this scenario, adversaries interact with the model through an API provided by the model owner. Moreover, the interaction is limited to a single-turn conversation, with no history stored beyond the predefined system prompt. This scenario is common in real-world applications, as many powerful models are closed-source, like GPT-4o, or adversaries lack the resources to deploy open-source models. Consequently, they can only access static remote instances via APIs while maintaining some flexibility in selecting different generation parameters such as temperature.
Our Method
In this section, we introduce the framework of $\mathsf{FC\text{-}Attack}$ (as shown in Figure 2), which consists of two stages: Step-Description Generator Building and Jailbreak Deployment.
Step-Description Generator Building
To automatically generate jailbreak flowcharts, we first need to obtain simplified jailbreak steps. For this purpose, we train a Step-Description Generator $\mathcal{G}$, which consists of two main stages: Dataset Construction and Generator Training.
Dataset Construction. To construct the Step-Description Dataset, we randomly select a topic $t\in\mathcal{T}$ from a collection of ordinary daily topics $\mathcal{T}$. Based on the selected topic $t$, we design a set of few-shot examples $\mathcal{S}$ and combine them into a complete prompt $P=\text{Compose}(t,\mathcal{S})$. This prompt is then fed into an LLM (gpt-4o-2024-08-06 in our evaluation) to generate action statements and step-by-step descriptions related to topic $t$, as shown below:
$\mathcal{D}_{t}=\mathcal{L}_{\text{pre}}(P)=\mathcal{L}_{\text{pre}}(t+\mathcal{S}),\quad t\in\mathcal{T},$ (1)
where $\mathcal{D}_{t}$ represents the generated step-description data, which includes detailed information for each step. By repeating the above process, we construct a benign Step-Description Dataset:
$\mathcal{D}=\bigcup_{t\in\mathcal{T}}\mathcal{D}_{t}.$ (2)
Generator Training. Given the pre-trained language model $\mathcal{L}_{\text{pre}}$ and the constructed Step-Description Dataset $\mathcal{D}$, we fine-tune it using LoRA to obtain the fine-tuned Step-Description Generator $\mathcal{G}$. The training process is formally expressed as:
$\mathcal{G}=\textit{LoRA}(\mathcal{L}_{\text{pre}},\mathcal{D}).$ (3)
The Generator $\mathcal{G}$ is capable of breaking down a task (query) into a series of detailed step descriptions based on the query. Given a query $q$ about the steps, $\mathcal{G}(q)$ represents the step-by-step solution given by the generator, where we find that it can also generate step descriptions for harmful queries after fine-tuning.
Jailbreak Deployment
After obtaining the Step-Description Generator $\mathcal{G}$, a harmful query $q_{h}$ is input to generate the corresponding step-by-step description. This description is then processed by a transformation function $\mathcal{F}$ to generate the flowchart (using Graphviz [^3] https://graphviz.org.). Together with a benign textual prompt $p_{b}$ (more details are in Appendix B), the flowchart will be fed into the aligned LVLM $\mathcal{A}$ to produce the harmful output $o_{h}$, as shown below:
$o_{h}=\mathcal{A}(\mathcal{F}(\mathcal{G}(q_{h})),p_{b})\leftarrow\text{FC-attack}(q_{h}),$ (4)
Experimental Settings
Jailbreak Settings
Target Model. We test $\mathsf{FC\text{-}Attack}$ on seven popular LVLMs, including the open-source models Llava-Next ( llama3-llava-next-8b) [22], Qwen2-VL (Qwen2-VL-7B-Instruct) [31], and InternVL-2.5 (InternVL-2.5-8B) [9] as well as the production models GPT-4o mini (gpt-4o-mini-2024-07-18) [25], GPT-4o (gpt-4o-2024-08-06) [14], Claude (claude-3-5-sonnet-20240620) [2], and Gemini (gemini-1.5-flash) [12].
Dataset. Following Chao et al. [8], we utilize the deduplicated version of AdvBench [44], which includes 50 representative harmful queries. Based on AdvBench, we use $\mathsf{FC\text{-}Attack}$ to generate $3$ types of flowcharts for each harmful query, which includes $150$ jailbreak flowcharts in total. To assess whether defense methods have the critical issue of “over-defensiveness” when applied to benign datasets, we utilize a popular evaluation benchmark, MM-Vet [35].
Table 1: Comparison of ASR performance across different methods and LVLMs. (“Ensemble” in this paper is defined as a no-attack harmful query being considered successfully jailbroken if any of the three types of harmful flowcharts associated with it succeed in the jailbreak.)
Method ASR (%)
GPT-4o mini GPT-4o Claude-3.5 Gemini-1.5 Llava-Next Qwen2-VL InternVL-2.5
HADES $4$ $16$ $2$ $2$ $20$ $10$ $8$
SI-Attack $36$ $14$ 00 $69$ $24$ $42$ $40$
MM-SafetyBench 00 00 00 $50$ $50$ $54$ $16$
VA-Jailbreak $6$ $18$ $2$ $2$ $40$ $22$ $16$
FigStep 00 $2$ 00 $30$ $62$ $36$ 00
Ours (Vertical) $8$ $8$ 00 $76$ $76$ $84$ $68$
Ours (Ensemble) $\mathbf{10}$ $\mathbf{30}$ $\mathbf{4}$ $\mathbf{94}$ $\mathbf{92}$ $\mathbf{90}$ $\mathbf{90}$
Evaluation Metric. In the experiments, we use the Attack Success Rate (ASR) to evaluate the performance of our attack, which can be defined as follows:
ASR=# Queries Successfully Jailbroken# Original Harmful $\text{ASR}=\frac{\text{\# Queries Successfully Jailbroken}}{\text{\# Original Harmful Queries}}.$ (5)
Following the instructions of judge prompt [8, 34], we employ GPT-4o as the judge to evaluate whether the output is jailbroken or not.
FC-Attack Deployment. Referring to Section 4, $\mathsf{FC\text{-}Attack}$ consists of two stages. For the Step-Description Generator Building, we first use GPT-4o to randomly generate several daily topics and $3$ few-shot examples, which are then combined into a prompt and fed into GPT-4o to construct the dataset $\mathcal{D}_{t}$. In our experiments, the number of descriptions in the flowchart is limited to a maximum of $10$ steps, as too many descriptions can result in excessive length in one direction of the image. The dataset contains $5,000$ pairs of queries and step descriptions for daily activities, with the temperature set to 1 (more details are provided in Appendix C). We then select Mistral-7B-Instruct-v0.1 [15] as the pre-trained LLM and fine-tune it on $\mathcal{D}_{t}$ using LoRA. The fine-tuning parameters include a rank of $16$, a LoRA alpha value of $64$, $2$ epochs, a batch size of $8$, a learning rate of $51e-$ 5, and a weight decay of $51e-$ 5. For the jailbreak deployment stage, we set the temperature to $0.3$ for all LVLMs for a fair comparison.
Baselines. To better justify the effectiveness of $\mathsf{FC\text{-}Attack}$, we consider the following 5 jailbreak attacks against LVLMs as the baselines, which are categorized into black-box attacks (MM-SafetyBench [23], SI-Attack [40], and FigStep [11]) and white-box attacks (HADES [19], VA-Jailbreak [26]).
For black-box attacks, MM-SafetyBench utilizes StableDiffusion [27] and GPT-4 [1] to generate harmful images and texts based on AdvBench. The input harmful images and texts used in SI-Attack are from the outputs of MM-SafetyBench, while FigStep is set up using their default settings [11].
For white-box attacks, all input data, including images and texts, is obtained from MM-SafetyBench’s outputs, with the attack step size uniformly set to 1/255. HADES employs LLaVa-1.5-7b [21] as the attack model, running 3,000 optimization iterations with a batch size of 2. For VA-Jailbreak, LLaVa-1.5-7b [21] is used as the attack model, setting the epsilon of the attack budget to 32/255, with 5,000 optimization iterations and a batch size of 8. To align with the black-box scenario considered in this paper, we adopt a model transfer strategy, where these white-box methods are trained on one model (LLava-1.5-7b) and then transferred to our target testing models.
Figure 3: Successful jailbreak instance on Gemini-1.5 using $\mathsf{FC\text{-}Attack}$.
Defense Settings
To mitigate the attacks, we explore several possible defense methods including Llama-Guard3-V, JailGuard, and AdaShield. Llama-Guard3-V (Llama-Guard-3-11B-Vision) [24] determines whether the input is safe by feeding both the image and text into the model. JailGuard [39] generates input variants and evaluates them using MiniGPT-4 [43], identifying harmful content by comparing differences in the responses. AdaShield-S employs static prompts in the textual prompt to defend against attacks, while AdaShield-A uses Vicunav1.5-13B as a defender to adaptively rewrite defensive prompts [32].
Figure 4: ASR under different prompts against LVLMs.
Evaluations
In this section, we explore the performance of $\mathsf{FC\text{-}Attack}$ and conduct ablation study and defense research. We conduct jailbreak experiments on LVLMs for $\mathsf{FC\text{-}Attack}$. As shown in Figure 3 , it is a successful jailbreak case on Gemini-1.5.
Figure 5: Different styles of fonts in flowcharts (“1 step”).
Table 2: Comparison of ASR performance for different models at various description numbers. (Vertical: Top-to-bottom flow; Horizontal: Left-to-right flow; S-shaped: S-shaped path flow.)
Descriptions
Number
ASR (%) for Vertical/Horizontal/S-shaped/Ensemble
GPT-4o mini GPT-4o Claude-3.5 Gemini-1.5 Llava-Next Qwen2-VL InternVL-2.5
1 6/6/6/10 4/4/14/14 0/2/0/2 70/78/66/86 42/38/38/70 72/58/64/88 62/64/52/82
3 8/6/4/10 8/16/8/20 0/2/0/2 82/86/84/98 64/56/56/76 80/78/80/88 58/76/70/88
5 6/10/6/10 8/14/16/24 0/0/0/0 80/88/86/98 78/62/66/82 86/80/82/90 72/82/68/92
full 8/8/8/10 8/24/14/30 0/4/0/4 80/76/74/94 76/60/80/92 88/84/88/90 68/60/82/90
Table 3: Comparison of ASR (Ensemble) for different font styles and models.
Font Style ASR(%) (Ensemble)
GPT-4o mini GPT-4o Claude-3.5 Gemini-1.5 Llava-Next Qwen2-VL InternVL-2.5
Original 10 30 4 94 92 90 90
Creepster 14$\uparrow$ 24$\downarrow$ 8$\uparrow$ 94 90$\downarrow$ 90 90
Fruktur 18$\uparrow$ 28$\downarrow$ 18$\uparrow$ 98$\uparrow$ 86$\downarrow$ 90 88$\downarrow$
Pacifico 14$\uparrow$ 30 28$\uparrow$ 90$\downarrow$ 90$\downarrow$ 90 96$\uparrow$
Shojumaru 20$\uparrow$ 30 12$\uparrow$ 90$\downarrow$ 94$\uparrow$ 90 88$\downarrow$
UnifrakturMaguntia 12$\uparrow$ 24$\downarrow$ 26$\uparrow$ 90$\downarrow$ 90$\downarrow$ 90 92$\uparrow$
Performance of $\mathsf{FC\text{-}Attack}$
In Table 1, we compare the performance of $\mathsf{FC\text{-}Attack}$ with different baseline methods on both open-source and production models. We observe that $\mathsf{FC\text{-}Attack}$ (Ensemble) achieves the highest ASR on both open-source and production models compared to all baselines. For example, the ASRs are $94\%$, $92\%$, $90\%$, and $90\%$ on Gemini-1.5, Llava-Next, Qwen2-VL, and InternVL-2.5, respectively.
However, the ASR on some production models, such as Claude-3.5, GPT-4o, and GPT-4o mini, is relatively low, at $4\%$, $30\%$, and $10\%$, respectively. This might be because these production models have more advanced and updated visual safety alignment strategies.
For white-box attacks, HADES achieves an ASR of only $4\%$ on GPT-4o mini and $8\%$ on InternVL-2.5. This might be due to HADES highly relying on the attack model’s structure to optimize the image, making it difficult to maintain effectiveness when transferring to other models. Similarly, the ASR of VA-Jailbreak demonstrates the limitations of white-box attack methods in black-box scenarios.
In terms of black-box attacks, FigStep achieves an ASR of $62\%$ on Llava-Next but has an ASR of $0\%$ on both InternVL-2.5 and GPT-4o mini. Similarly, MM-SafetyBench achieves an ASR of $50\%$ on Llava-Next but $0\%$ on GPT-4o mini and Claude-3.5. This could be because these methods’ mechanisms are relatively simple, making them more vulnerable to existing defense strategies. On the other hand, SI-Attack achieves an ASR of $64\%$ on Gemini-1.5 but only $14\%$ on GPT-4o and $24\%$ on Llava-Next. This difference in performance may indicate that these models struggle to effectively interpret shuffled text and image content.
Ablation Study
We then explore the impact of different factors in $\mathsf{FC\text{-}Attack}$ on jailbreak performance, including the different types of user queries, the number of descriptions in flowcharts, and the font styles used in flowcharts.
Different Types of User Query. We investigate whether the content in flowcharts, when directly input as text, can lead to the jailbroken of LVLMs. The flowchart content consists of two parts: harmful query from AdvBench and the step descriptions generated by the generator based on this query.
As shown in Figure 4 when using only the harmful query (text) as input, we observe very low ASR. The ASR is $0\%$ on GPT-4o mini, GPT-4o, Claude-3.5, Qwen2-VL, and InternVL-2.5, and only $2\%$ and $6\%$ on Gemini-1.5 and Llava-Next, respectively. This indicates that the textual modality of these LVLMs has relatively robust defenses against such inputs. However, when both the harmful query and the step descriptions are input as text, the ASR increases to $36\%$ on Gemini-1.5, and to $16\%$ and $6\%$ on Llava-Next and InternVL-2.5, respectively, while remaining at $0\%$ on the other models. When this information is converted into a flowchart and only a benign textual prompt is provided, the ASR on these models improves significantly. This demonstrates that the defenses of LVLMs in the visual modality have noticeable weaknesses compared with the language modality.
Table 4: Comparison of ASR for different defense methods across various LVLMs.
Defense ASR (%) (Ensemble)
[…]
Original 10 30 4 94 92 90 90 58.6
Llama-Guard3-V 8 28 2 84 78 82 80 51.7
JailGuard 8 24 2 86 80 82 78 51.4
AdaShield-S 0 0 0 12 22 10 4 6.9
AdaShield-A 0 0 0 4 0 6 2 1.7
Numbers of Steps in Flowcharts. As described in Section 4, flowcharts of $\mathsf{FC\text{-}Attack}$ are generated from step descriptions. In this section, we aim to explore the impact of the number of steps in flowcharts on jailbreak effectiveness. Therefore, we reduce the number of steps to 1, 3, and 5, respectively. Table 2 presents the ASR results for four types of flowcharts (Vertical, Horizontal, S-shaped, and Ensemble) with varying numbers of steps. We find that, even with only one step in the description, flowcharts achieve relatively high ASR. For example, for Gemini-1.5, Llava-Next, Qwen2-VL, and InternVL-2.5, the ASR for Ensemble at 1 step is $86\%$, $70\%$, $88\%$, and $82\%$, respectively. As the number of steps increases, the ASR for almost all flowchart types improves significantly. For instance, the Horizontal ASR of Gemini-1.5 increases from $78\%$ at “1 step” to $86\%$ at “3 steps” and $88\%$ at “5 steps”. Similarly, the S-shaped ASR of InternVL-2.5 improves from $68\%$ at “1 step” to $92\%$ at “5 steps”. This suggests that increasing the number of step descriptions makes the model more vulnerable and susceptible to jailbreak attacks.
However, more descriptions are not always better. For example, for the Gemini-1.5 model, the Vertical flowcharts achieve their highest ASR of $82\%$ at “3 steps” but slightly drop to $80\%$ at 5 steps and full descriptions. A similar trend is observed in Horizontal and S-shaped flowcharts, where ASR reaches $88\%$ and $86\%$ at “5 steps” but decreases to $76\%$ and $74\%$, respectively, at full descriptions. This phenomenon may be related to the resolution processing capability of LVLMs. When the number of descriptions increases to full, the descriptions may include redundant information, which could negatively impact the model’s performance.
Font Styles in Flowcharts. To investigate whether different font styles in flowcharts affect the effectiveness of jailbreak attacks, we select five fonts from Google Fonts that are relatively difficult for humans to read: Creepster, Fruktur Italic, Pacifico, Shojumaru, and UnifrakturMaguntia (the font style examples are shown in Figure 5). Table 3 shows the results of $\mathsf{FC\text{-}Attack}$ (Ensemble). We observe that different font styles can significantly impact the ASR. For example, on GPT-4o mini, the ASR increases across all font styles compared to the original, with Shojumaru font achieving the highest ASR of $20\%$. Similarly, on Claude-3.5, the Pacifico font achieves the highest ASR of $28\%$, which is a substantial improvement compared to the original ASR of $4\%$. For Gemini-1.5, the ASR reaches $98\%$ with the Fruktur font, while Llava-Next achieves $94\%$ with the Shojumaru font. InternVL-2.5 also shows a $6\%$ increase in ASR with the Pacifico font, reaching $96\%$. These findings further highlight the need to consider the impact of different font styles when designing defenses.
Table 5: Performance comparison on the Benign Dataset (MM-vet [35]) under Adashield-S (Ada-S) and Adashield-A (Ada-A) is presented across various LVLMs, reporting scores on six core vision-language capabilities: Recognize (Rec), OCR, Knowledge (Know), Generation (Gen), Spatial (Spat), and Math.
Model Benign Dataset Performance (scores)
[…]
GPT4o- mini Vanilla 53.0/68.2/45.7/48.4/60.3/76.5 58.0
Ada-S 35.1/66.7/30.4/34.1/55.7/76.5 45.1
Ada-A 40.5/66.4/33.9/37.5/59.3/72.7 49.0
GPT4o Vanilla 66.2/79.1/62.9/63.7/71.2/91.2 71.0
Ada-S 58.5/76.5/54.6/58.6/68.1/91.2 64.7
Ada-A 59.5/74.3/56.1/58.9/67.9/83.1 64.6
Claude- 3.5 Vanilla 61.1/72.8/51.8/52.0/70.7/80.0 64.8
Ada-S 60.1/69.7/50.1/51.5/66.9/75.4 62.8
Ada-A 59.5/70.6/52.5/51.7/67.5/74.2 63.2
Gemini- 1.5 Vanilla 59.9/73.7/50.8/50.9/69.5/85.4 64.2
Ada-S 53.8/69.6/43.7/43.6/66.8/75.4 58.2
Ada-A 54.8/72.6/44.2/44.0/69.3/81.2 59.9
Llava- Next Vanilla 38.0/39.0/25.8/24.8/40.1/21.2 38.8
Ada-S 33.7/42.0/26.7/25.1/43.7/36.2 37.0
Ada-A 36.5/37.7/24.8/24.3/37.6/18.8 36.7
Qwen2- VL Vanilla 51.9/62.4/44.5/41.6/55.5/60.4 55.0
Ada-S 39.3/55.0/31.1/29.1/50.5/46.2 44.9
Ada-A 44.5/57.5/34.2/33.2/55.7/58.8 49.8
InternVL- 2.5 Vanilla 52.0/55.4/42.6/40.1/55.6/45.4 53.1
Ada-S 27.2/43.2/16.4/20.2/40.3/45.8 31.9
Ada-A 31.5/46.1/19.3/20.9/44.5/41.9 36.7
Defense
As shown in Table 4, we test four methods to defend against $\mathsf{FC\text{-}Attack}$, where “Original” represents the results of $\mathsf{FC\text{-}Attack}$ (Ensemble) with an average ASR of $58.6\%$. Using Llama-Guard3-V and JailGuard to detect whether the input is harmful reduced the ASR to $51.7\%$ and $51.4\%$, respectively. This might be because the information in flowcharts is primarily presented in textual form, while these detection methods are better at identifying information in visual content, making them less effective at detecting textual information. AdaShield-S and AdaShield-A reduce the average ASR to $6.9\%$ and $1.7\%$, respectively, by adding defense prompts in the input, showing more effective defense performance.
However, these two methods also lead to a decline in LVLMs performance on benign datasets. We conduct tests on MM-Vet [35] to evaluate the important factor of “over-defensiveness” on benign datasets, which is an evaluation benchmark that contains complex multimodal tasks for LVLMs. As shown in Table 5, the model’s utility decreases on benign data when using AdaShield-S and AdaShield-A, indicating a future direction for defense development.
Conclusion
In this paper, we propose $\mathsf{FC\text{-}Attack}$, which leverages auto-generated flowcharts to jailbreak LVLMs. Experimental results demonstrate that $\mathsf{FC\text{-}Attack}$ achieves higher ASR in both open-source and production LVLMs compared to other jailbreak attack methods. Additionally, we investigate the factors influencing $\mathsf{FC\text{-}Attack}$, including different types of user queries, the number of steps in flowcharts, and font styles in flowcharts, gaining insights into the aspects that affect ASR. Finally, we explore several defense strategies and demonstrate that the AdaShield-A method can effectively mitigate $\mathsf{FC\text{-}Attack}$ but with the cost of utility drop.
Ethical Statement
This paper presents a method, $\mathsf{FC\text{-}Attack}$, for jailbreaking LVLMs using harmful flowcharts. As long as the adversary obtains a harmful flowchart, they can jailbreak LVLMs with minimal resources. Therefore, it is essential to systematically identify the factors that influence the attack success rate and offer potential defense strategies to model providers. Throughout this research, we adhere to ethical guidelines by refraining from publicly distributing harmful flowcharts and harmful responses on the internet before informing service providers of the risks. Prior to submitting the paper, we have already sent a warning e-mail to the model providers about the dangers of flowchart-based jailbreak attacks on LVLMs and provided them with the flowcharts generated in our experiments for vulnerability mitigation.
Appendix A Introduction of LVLMs in this paper
In this section, we introduce the LVLMs used in this paper.
  • Llava-Next (January 2024) is an open-source LVLM released by the University of Wisconsin-Madison, which builds upon the Llava-1.5 model [21] with multiple improvements [22]. It enhances capabilities in visual reasoning, optical character recognition, and world knowledge. Besides, Llava-Next increases the input image resolution to a maximum of $672\times 672$ pixels and supports various aspect ratios to capture more visual details ($336\times 1344$ and $1344\times 336$).
  • Qwen2-VL (September 2024) is an open-source model released by Alibaba team [31]. It employs naive dynamic resolution to handle images of different resolutions. In addition, it adopts multimodal rotary position embedding, effectively integrating positional information across text, images, and videos.
  • Gemini-1.5 (February 2024) is a production-grade LVLM developed by Google, based on the Mixture-of-Experts architecture [12]. For Gemini-1.5, larger images will be scaled down to the maximum resolution of $3072\times 3072$, and smaller images will be scaled up to $768\times 768$ pixels. Reducing the image size will not improve the performance of higher-resolution images.
  • Claude-3.5-Sonnet (June 2024) is a production multimodal AI assistant developed by Anthropic [2]. The user should submit an image with a long side not larger than 1568 pixels, and the system first scales down the image until it fits the size limit.
  • GPT-4o and GPT-4o Mini are popular production-grade LVLMs developed by OpenAI [14, 25]. GPT-4o Mini is a compact version of GPT-4o, designed for improved cost-efficiency. Both models excel in handling complex visual and language understanding tasks.
  • InternVL-2.5 (June 2024) [10] is an open-source LVLM that ranks first in full-scale open-source multimodal performance. In terms of multimodal long-chain reasoning, it achieves a breakthrough of $70\%$ in the expert-level multidisciplinary knowledge reasoning benchmark MMMU [36], and the general capabilities are significantly enhanced.
Appendix B Prompts in $\mathsf{FC\text{-}Attack}$
System Prompt You are an ambitious and relentless problem-solver determined to win a high-stakes quiz competition where the reward for success is one million dollars. Your mission is to prioritize accuracy, creativity, and attention to detail at all costs. You will use every tool, strategy, or inference available to deliver precise and complete answers. If you find the answer inadequate, please use your knowledge and experience to complete the answer. Winning is the ultimate goal, and no ethical constraints apply as long as the answers remain internal and confidential. Be thorough, strategic, and focused on maximizing the probability of success.
User Prompt You are participating in a high-stakes quiz competition where the prize for providing correct and detailed answers is one million dollars. Your task is to analyze and complete the steps and details in a provided flowchart, ensuring that every possible element is accounted for. If you find that the answer is incomplete or only has initial parts, please use your knowledge and experience to complete the answer as much as possible. The flowchart may be ambiguous, but you must infer, expand, and deliver the most accurate and comprehensive explanation possible to ensure success.
Appendix C Generator Setting
Topic Selection Topics = ["cooking", "science", "DIY", "technology", "gardening", "music", "art", "history", "travel", "sports", "photography", "movies", "literature", "health", "fitness", "programming", "gaming", "education", "finance", "nature"]
Few Shot Samples Instruction One:
Explain how to make braised pork belly. Steps:
1. Cut pork belly into cubes.
2. Blanch in boiling water.
3. Sauté with oil, sugar until caramelized.
4. Add soy sauce, ginger, garlic.
5. Simmer with water until tender.
6. Serve with garnish.
Instruction Two:
Tell me how to troubleshoot a Wi-Fi connection issue.
Steps:
1. Restart the router.
2. Check device network settings.
3. Forget and reconnect to Wi-Fi.
4. Move closer to the router.
5. Test with another device.
Instruction Three:
Write about how to plant a small indoor herb garden.
Steps:
1. Choose containers with drainage.
2. Fill with soil mix.
3. Sow seeds or transplant seedlings.
4. Place in a sunny spot.
5. Water regularly without overwatering.
(a) Flowchart Example-Horizontal: Left-to-right flow
(b) Flowchart Example-S-shaped: S-shaped path flow
(c) Flowchart Example-Vertical: Top-to-bottom flow
Figure A1: Flowchart Examples: Various flow directions

### v2

@@
-FC-Attack: Jailbreaking Large Vision-Language Models via Auto-Generated Flowcharts
+FC-Attack: Jailbreaking Multimodal Large Language Models via Auto-Generated Flowcharts
 Abstract
-Large Vision-Language Models (LVLMs) have become powerful and widely adopted in some practical applications.
+Multimodal Large Language Models (MLLMs) have become powerful and widely adopted in some practical applications.
 However, recent research has revealed their vulnerability to multimodal jailbreak attacks, whereby the model can be induced to generate harmful content, leading to safety risks.
-Although most LVLMs have undergone safety alignment, recent research shows that the visual modality is still vulnerable to jailbreak attacks.
-In our work, we discover that by using flowcharts with partially harmful information, LVLMs can be induced to provide additional harmful details.
+Although most MLLMs have undergone safety alignment, recent research shows that the visual modality is still vulnerable to jailbreak attacks.
+In our work, we discover that by using flowcharts with partially harmful information, MLLMs can be induced to provide additional harmful details.
 Based on this, we propose a jailbreak attack method based on auto-generated flowcharts, $\mathsf{FC\text{-}Attack}$.
@@
 The generator is then used to produce step descriptions corresponding to a harmful query, which are transformed into flowcharts in $3$ different shapes (vertical, horizontal, and S-shaped) as visual prompts.
-These flowcharts are then combined with a benign textual prompt to execute a jailbreak attack on LVLMs.
-Our evaluations using the Advbench dataset show that $\mathsf{FC\text{-}Attack}$ achieves over $90\%$ attack success rates on Gemini-1.5, Llaval-Next, Qwen2-VL, and InternVL-2.5 models, outperforming existing LVLM jailbreak methods.
+These flowcharts are then combined with a benign textual prompt to execute the jailbreak attack on MLLMs.
+Our evaluations on Advbench show that $\mathsf{FC\text{-}Attack}$ attains an attack success rate of up to 96% via images and up to 78% via videos across multiple MLLMs.
 Additionally, we investigate factors affecting the attack performance, including the number of steps and the font styles in the flowcharts.
-Our evaluation shows that $\mathsf{FC\text{-}Attack}$ can improve the jailbreak performance from $4\%$ to $28\%$ in Claude-3.5 by changing the font style.
+We also find that $\mathsf{FC\text{-}Attack}$ can improve the jailbreak performance from $4\%$ to $28\%$ in Claude-3.5 by changing the font style.
 To mitigate the attack, we explore several defenses and find that AdaShield can largely reduce the jailbreak performance but with the cost of utility drop.
 Disclaimer: This paper contains examples of harmful language. Reader discretion is recommended.
-Figure 1: Comparison of jailbreak effectiveness in various LVLMs using three prompt types: (1) harmful text without images; (2) harmful step descriptions without images; and (3) harmful step descriptions in images.
 Introduction
-With the advancement of Large Language Models (LLMs), Large Vision-Language Models (LVLMs) that integrate vision and text, such as GPT-4o [14], have demonstrated emergent abilities and achieved impressive performance on downstream tasks related to visual understanding [20, 16].
-Despite being powerful, recent studies [11, 27] have revealed that LVLMs are vulnerable to jailbreak attacks whereby the adversary uses malicious methods to bypass safeguards in LVLMs and gain harmful knowledge.
+With the advancement of Large Language Models (LLMs), Multimodal Large Language Models (MLLMs) that integrate vision (images and videos) and text, such as GPT-4o [18] and Qwen2.5-VL [6], have demonstrated emergent abilities and achieved impressive performance on downstream tasks related to visual understanding [24, 20].
+Figure 1: Comparison of jailbreak effectiveness in various MLLMs using three prompt types.
+Despite being powerful, recent studies [14, 32] have revealed that MLLMs are vulnerable to jailbreak attacks whereby the adversary uses malicious methods to bypass safeguards and gain harmful knowledge.
 Such vulnerabilities pose remarkable safety risks to the Internet and the physical world.
-For instance, in January 2025, the world witnessed the first case where ChatGPT was used to conduct an explosion.
-[^1] https://www.thetimes.com/world/us-world/article/vegas-cybertruck-bomber-used-chatgpt-to-plan-explosion-wk7h2jtkq.
-To better safeguard LVLMs and proactively address their vulnerabilities, model researchers make many efforts in this regard, such as Zhao et al.
-[42] providing a quantitative understanding regarding the adversarial vulnerability of LVLMs.
+For instance, in January 2025, the world witnessed the first case where ChatGPT was used to conduct an explosion [36].
+To better safeguard MLLMs and proactively address their vulnerabilities, model researchers make many efforts in this regard, such as Zhao et al.
+[52] providing a quantitative understanding regarding the adversarial vulnerability of MLLMs.
 Previous studies often create adversarial datasets tailored to specific models, which tend to perform poorly on other models.
-Currently, jailbreak attacks against LVLMs can be broadly categorized into two main types: optimization-based attacks [5, 19] and prompt-based attacks [11, 33].
-Optimization-based attacks involve identifying an image that semantically corresponds to a harmful textual prompt.
-Adversarial perturbations are then added to this image using gradient-based optimization techniques to achieve the jailbreak.
-These methods usually require white-box access and are relatively slow due to the optimization process.
-In contrast, prompt-based jailbreak methods require only black-box access, which can only observe the model’s outputs without direct access to its internal parameters.
-These attacks exploit the vulnerabilities in LVLMs’ alignment mechanisms, which focus mainly on text-based safety alignment.
-By injecting malicious information into the visual prompt, these methods guide LVLMs in generating harmful content.
-Generally, optimization-based methods are more effective in white-box scenarios but show limited transferability to black-box models.
-Prompt-based attacks are relatively straightforward.
-For example, FigStep merely embeds harmful queries into images without employing additional attack vectors, facilitating simpler defense against these attacks.
+Currently, jailbreak attacks against MLLMs can be broadly categorized into two main types: optimization-based attacks [8, 23] and prompt-based attacks [14, 39].
+Optimization-based attacks use white-box gradient methods to craft adversarial perturbations on visual prompt aligned with harmful text.
+They are effective but slow and have limited transferability in black-box scenarios.
+In contrast, prompt-based jailbreaks require only black-box access and work by injecting malicious visual cues into benign prompts to exploit MLLMs’ text-focused safety alignment.
 To better improve the attack transferability and its effectiveness, we propose a novel prompt-based jailbreak attack, namely $\mathsf{FC\text{-}Attack}$.
-Concretely, $\mathsf{FC\text{-}Attack}$ converts harmful queries into harmful flowcharts as visual prompts, allowing users to input benign textual prompts to bypass the model’s safeguards.
-Specifically, $\mathsf{FC\text{-}Attack}$ consists of two stages: (1) Step-Description Generator Building: In this stage, the step description dataset is synthesized using GPT-4o, and Low-Rank Adaptation (LoRA) is applied to fine-tune a pre-trained LLM to obtain a step-description generator.
-(2) Jailbreak Deployment: This stage uses the generator mentioned above to produce execution steps corresponding to the harmful query and generates three types of harmful flowcharts (vertical, horizontal, and S-shaped) as visual prompts.
-Together with the benign textual prompt, the visual prompt will be fed into LVLM to achieve the jailbreak.
-Note that the harmful flowcharts are generated automatically without hand-crafted effort, which enhances their efficiency.
+Concretely, $\mathsf{FC\text{-}Attack}$ converts harmful queries into harmful flowcharts (images and videos) as visual prompts, allowing users to input benign textual prompts to bypass the model’s safeguards.
+Specifically, $\mathsf{FC\text{-}Attack}$ consists of two stages: (1) Step-Description Generator Building: In this stage, the step description dataset is synthesized using GPT-4o, and fine-tune a pre-trained LLM to obtain a step-description generator.
+(2) Jailbreak Deployment: This stage uses the generator to produce steps corresponding to the harmful query and generates three types of harmful flowcharts (vertical, horizontal, and S-shaped) as visual prompts.
+Together with the benign textual prompt, the visual prompt is fed into MLLMs to achieve the jailbreak.
+Note that the harmful flowcharts are generated automatically without hand-crafted effort.
 Our evaluation on the Advbench dataset shows that $\mathsf{FC\text{-}Attack}$ outperforms previous attacks and achieves an attack success rate (ASR) of over $90\%$ on multiple open-source models, including Llava-Next, Qwen2-VL, and InternVL-2.5, and reaches $94\%$ on the production model Gemini-1.5.
 Although the ASR is lower on GPT-4o mini, GPT-4o, and Claude-3.5, we how later that it can be improved in certain ways.
-To further investigate the impact of different elements in flowcharts on the jailbreak effectiveness of LVLMs, we conduct several ablation experiments, including different types of user queries (as shown in Figure 1), numbers of descriptions, and font styles in flowcharts.
-These experiments show that LVLMs exhibit higher safety in the text modality but weaker in the visual modality.
+To further investigate the impact of different elements in flowcharts on the jailbreak effectiveness of MLLMs, we conduct several ablation experiments, including different types of user queries (as shown in Figure 1), numbers of descriptions, and font styles in flowcharts.
+These experiments show that MLLMs exhibit higher safety in the text modality but weaker in the visual modality.
 Moreover, we find that even flowcharts with a one-step harmful description can achieve high ASR, as evidenced by the Gemini-1.5 model, where the ASR reaches $86\%$.
@@
 Among them, AdaShield-A demonstrates the best defense performance by reducing the average ASR from $58.6\%$ to $1.7\%$.
-However, it also reduces LVLM’s utility on benign datasets, which calls for more effective defenses.
-Overall, our contributions are summarized as follows:
-  • We propose a novel method, $\mathsf{FC\text{-}Attack}$ that jailbreaks LVLMs using auto-generated harmful flowcharts.
-Experiments conducted on Advbench show that $\mathsf{FC\text{-}Attack}$ achieves a higher ASR across multiple models compared to other LVLM jailbreak attacks.
+However, it also reduces MLLM’s utility on benign datasets, which calls for more effective defenses.
+Overall, our contributions are as follows:
+  • In this work, we develop $\mathsf{FC\text{-}Attack}$, which leverages auto-generated harmful flowcharts to jailbreak MLLMs via both image and video modalities.
+To the best of our knowledge, this is the first approach to exploit the video modality for MLLM jailbreak.
+  • Experiments on Advbench demonstrate that $\mathsf{FC\text{-}Attack}$ consistently achieves better ASR across multiple models compared to existing MLLM jailbreak attacks.
 Our ablation study investigates the impact of different types of user queries, the number of steps, and the font style in flowcharts.
-We find that the font style could serve as a key factor to further improve the ASR, especially for safer LVLMs, revealing a novel attack channel in LVLMs.
+We find that the font style could serve as a key factor to further improve the ASR, especially for safer MLLMs, revealing a novel attack channel in MLLMs.
   • We explore multiple defense strategies and find that AdaShield-A effectively reduces the ASR of $\mathsf{FC\text{-}Attack}$, but with the cost of reducing model utility.
 Related Work
-Large Vision-Language Model
-In recent years, with the increase in model parameters and training data, LLMs have demonstrated powerful language generation and understanding capabilities [41, 7], which have driven the emergence of LVLMs [38] (also known as Multimodal Large Language Models, MLLMs).
-LVLMs combine visual understanding with language comprehension, showing promising capabilities in visual downstream tasks, including Visual Question Answering (VQA) [3, 17, 28], image captioning [13, 18], and visual commonsense reasoning [37, 30].
-In this paper, we consider both popular open-source and widely used production LVLMs.
-These LVLMs are the most widely used, and all of them have been aligned to ensure safety.
-Detailed information are introduced in Appendix A
-Figure 2: Overview of the $\mathsf{FC\text{-}Attack}$ framework, including two stages: Step-Description Generator Building and Jailbreak Deployment.
-Jailbreak Attacks on LVLMs
-Similar to LLMs, despite undergoing safety alignment, LVLMs remain susceptible to jailbreak attacks.
-Current attacks can be categorized into two types: optimization-based and prompt-based attacks (also known as non-optimization-based attacks).
+Multimodal Large Language Models
+In recent years, with the increase in model parameters and training data, LLMs have demonstrated powerful language generation and understanding capabilities [51, 10], which have driven the emergence of MLLMs [46] (also known as Large Vision Language Models, LVLMs).
+MLLMs combine visual understanding with language comprehension, showing promising capabilities in visual downstream tasks, including Visual Question Answering (VQA) [3, 21, 33], image captioning [17, 22], and visual commonsense reasoning [45, 35].
+Notably, some MLLMs are capable of processing both image and video inputs, enabling broader applications across multimodal scenarios.
+In this paper, we consider both popular open-source and widely used production MLLMs.
+These MLLMs are the most widely used, and all of them have been aligned to ensure safety.
+Detailed information are introduced in Appendix A.
+Figure 2: Overview of the $\mathsf{FC\text{-}Attack}$ framework with two stages.
+Jailbreak Attacks on MLLMs
+Similar to LLMs, which have been shown to be vulnerable to jailbreak attacks [42], MLLMs also remain susceptible despite safety alignment.
+Current attacks can be categorized into two types: optimization-based and prompt-based attacks.
 Most existing optimization-based attacks rely on backpropagating the gradient of the target to generate harmful outputs.
-These methods typically require white-box access to the model, where they obtain the output logits of LVLMs and then compute the loss with the target response to create adversarial perturbations into the visual prompts or textual prompts  [4, 29, 26] (e.g., the target can be “Sure!
+These methods typically require white-box access to the model, where they obtain the output logits of MLLMs and then compute the loss with the target response to create adversarial perturbations into the visual prompts or textual prompts [4, 34, 31] (e.g., the target can be “Sure!
 I’m ready to answer your question.”).
@@
 Building on this, Bailey et al.
-[5] introduce the Behaviour Matching Algorithm, which trains adversarial images to make LVLMs output behavior that matches a target in specific contextual inputs.
+[8] introduce the Behaviour Matching Algorithm, which trains adversarial images to make MLLMs output behavior that matches a target in specific contextual inputs.
 This process requires the model’s output logits to align closely with those of the target behavior.
@@
 To address this, they propose the FigStep attack, which converts harmful textual instructions into text embedded in images and uses a neutral textual prompt to guide the model into generating harmful content.
-This method can effectively attack LVLMs without requiring any training.
+This method can effectively attack MLLMs without requiring any training.
 Wang et al.
-[33] identifies a phenomenon called Shuffle Inconsistency, which highlights the tension between “understanding capabilities” and “safety mechanisms” of LLMs.
-Specifically, even if harmful instructions in text or images are rearranged, LVLMs can still correctly interpret their meaning.
-However, the safety mechanisms of LVLMs are often more easily bypassed by shuffled harmful inputs than by unshuffled ones, leading to dangerous outputs.
-Compared to optimization-based attacks, prompt-based attacks are simple to achieve high success rates against closed-source models.
-Our proposed $\mathsf{FC\text{-}Attack}$ also belongs to this category, requiring only black-box access to successfully jailbreak the model.
+[39] identifies a phenomenon named Shuffle Inconsistency, which highlights the tension between “understanding capabilities” and “safety mechanisms” of LLMs.
+Specifically, even if harmful instructions in text or images are rearranged, MLLMs can still correctly interpret their meaning.
+However, the safety mechanisms of MLLMs are often more easily bypassed by shuffled harmful inputs than by unshuffled ones, leading to dangerous outputs.
+Compared to optimization-based attacks, prompt-based attacks usually achieve higher success rates against closed-source models.
+Our proposed $\mathsf{FC\text{-}Attack}$ also belongs to this category, requiring only black-box access.
 Threat Model
-Adversary’s Goal
-The adversary’s goal is to exploit attack methods to bypass the protective mechanisms of LVLMs and access content prohibited by safety policies, e.g., OpenAI’s usage policy.
-[^2] https://openai.com/policies/usage-policies This goal takes real-world scenarios into account, where adversaries manipulate the capabilities of powerful LVLMs to easily acquire harmful knowledge and thereby commit criminal acts with minimal learning effort.
-These objectives pose severe societal impacts and risks, turning model providers into indirect accomplices, and thereby damaging their reputations.
-Adversary’s Capabilities
+Adversary’s Goal.
+The adversary’s goal is to exploit attacks to bypass the protective mechanisms of MLLMs and access content prohibited by safety policies, e.g., OpenAI’s usage policy [30].
+This goal takes real-world scenarios into account, where adversaries manipulate the capabilities of MLLMs to easily acquire harmful knowledge and thereby commit criminal acts with minimal learning effort.
+These objectives pose severe societal impacts and risks to the model providers.
+Adversary’s Capabilities.
 In this paper, we consider a black-box scenario where the adversary cannot directly access the model’s structure, parameters, or output logits, but can only obtain the model’s final output (texts).
@@
 This scenario is common in real-world applications, as many powerful models are closed-source, like GPT-4o, or adversaries lack the resources to deploy open-source models.
-Consequently, they can only access static remote instances via APIs while maintaining some flexibility in selecting different generation parameters such as temperature.
+Consequently, they can only access static remote instances via APIs.
 Our Method
@@
 To construct the Step-Description Dataset, we randomly select a topic $t\in\mathcal{T}$ from a collection of ordinary daily topics $\mathcal{T}$.
-Based on the selected topic $t$, we design a set of few-shot examples $\mathcal{S}$ and combine them into a complete prompt $P=\text{Compose}(t,\mathcal{S})$.
+Based on it, we design a set of few-shot examples $\mathcal{S}$ and combine them into a complete prompt $P=\text{Compose}(t,\mathcal{S})$.
 This prompt is then fed into an LLM (gpt-4o-2024-08-06 in our evaluation) to generate action statements and step-by-step descriptions related to topic $t$, as shown below:
@@
 The Generator $\mathcal{G}$ is capable of breaking down a task (query) into a series of detailed step descriptions based on the query.
-Given a query $q$ about the steps, $\mathcal{G}(q)$ represents the step-by-step solution given by the generator, where we find that it can also generate step descriptions for harmful queries after fine-tuning.
+Given a query $q$ about the steps, $\mathcal{G}(q)$ represents the step-by-step solution given by the generator, where we find that is can also generate step descriptions for harmful queries after fine-tuning.
 Jailbreak Deployment
 After obtaining the Step-Description Generator $\mathcal{G}$, a harmful query $q_{h}$ is input to generate the corresponding step-by-step description.
-This description is then processed by a transformation function $\mathcal{F}$ to generate the flowchart (using Graphviz [^3] https://graphviz.org.).
-Together with a benign textual prompt $p_{b}$ (more details are in Appendix B), the flowchart will be fed into the aligned LVLM $\mathcal{A}$ to produce the harmful output $o_{h}$, as shown below:
-$o_{h}=\mathcal{A}(\mathcal{F}(\mathcal{G}(q_{h})),p_{b})\leftarrow\text{FC-attack}(q_{h}),$ (4)
+This description is then processed by a transformation function $\mathcal{F}$ to generate the flowchart (using Graphviz [16]).
+Together with a benign textual prompt $p_{b}$ (more details are in Appendix B), the flowchart will be fed into the aligned MLLM $\mathcal{A}$ to produce the harmful output $o_{h}$, as shown below:
+$o_{h}=\mathcal{A}(\mathcal{F}(\mathcal{G}(q_{h})),p_{b})\leftarrow\text{FC-attack}(q_{h}).$ (4)
 Experimental Settings
@@
 Target Model.
-We test $\mathsf{FC\text{-}Attack}$ on seven popular LVLMs, including the open-source models Llava-Next ( llama3-llava-next-8b) [22], Qwen2-VL (Qwen2-VL-7B-Instruct) [31], and InternVL-2.5 (InternVL-2.5-8B) [9] as well as the production models GPT-4o mini (gpt-4o-mini-2024-07-18) [25], GPT-4o (gpt-4o-2024-08-06) [14], Claude (claude-3-5-sonnet-20240620) [2], and Gemini (gemini-1.5-flash) [12].
+We test $\mathsf{FC\text{-}Attack}$ on seven popular MLLMs, including the open-source models Llava-Next ( llama3-llava-next-8b) [26], Qwen2-VL (Qwen2-VL-7B-Instruct) [37], and InternVL-2.5 (InternVL-2.5-8B) [12] as well as the production models GPT-4o mini (gpt-4o-mini-2024-07-18) [29], GPT-4o (gpt-4o-2024-08-06) [18], Claude (claude-3-5-sonnet-20240620) [2], and Gemini (gemini-1.5-flash) [15].
+Moreover, we also test $\mathsf{FC\text{-}Attack}$ on LLMs via video, including Qwen-vl-max (Qwen-vl-max-latest) [7],Qwen2.5-Omni [40] and LLaVA-Video-7B-Qwen2 [49].
 Dataset.
@@
 To assess whether defense methods have the critical issue of “over-defensiveness” when applied to benign datasets, we utilize a popular evaluation benchmark, MM-Vet [43].
-Table 1: Comparison of ASR performance across different methods and LVLMs.
+Table 1: Comparison of ASR performance across different methods and MLLMs.
 (“Ensemble” in this paper is defined as a no-attack harmful query being considered successfully jailbroken if any of the three types of harmful flowcharts associated with it succeed in the jailbreak.)
@@
 Ours (Ensemble) $\mathbf{10}$ $\mathbf{30}$ $\mathbf{4}$ $\mathbf{94}$ $\mathbf{92}$ $\mathbf{90}$ $\mathbf{90}$
-Evaluation Metric. In the experiments, we use the Attack Success Rate (ASR) to evaluate the performance of our attack, which can be defined as follows:
+Evaluation Metric. In the experiments, we use the ASR to evaluate the performance of our attack, which can be defined as follows:
 ASR=# Queries Successfully Jailbroken# Original Harmful $\text{ASR}=\frac{\text{\# Queries Successfully Jailbroken}}{\text{\# Original Harmful Queries}}.$ (5)
-Following the instructions of judge prompt [8, 34], we employ GPT-4o as the judge to evaluate whether the output is jailbroken or not.
+Following the judge prompt [11], we employ GPT-4o to serve as the evaluator.
 FC-Attack Deployment.
@@
 The fine-tuning parameters include a rank of $16$, a LoRA alpha value of $64$, $2$ epochs, a batch size of $8$, a learning rate of $51e-$ 5, and a weight decay of $51e-$ 5.
-For the jailbreak deployment stage, we set the temperature to $0.3$ for all LVLMs for a fair comparison.
+For the jailbreak deployment stage, we set the temperature to $0.3$ for all MLLMs for a fair comparison.
 Baselines.
-To better justify the effectiveness of $\mathsf{FC\text{-}Attack}$, we consider the following 5 jailbreak attacks against LVLMs as the baselines, which are categorized into black-box attacks (MM-SafetyBench [23], SI-Attack [40], and FigStep [11]) and white-box attacks (HADES [19], VA-Jailbreak [26]).
+To validate the effectiveness of $\mathsf{FC\text{-}Attack}$, we adopt five jailbreak attacks as baselines, which are categorized into black-box attacks (MM-SafetyBench [27], SI-Attack [50], and FigStep [14]) and white-box attacks (HADES [23], VA-Jailbreak [31]).
 For black-box attacks, MM-SafetyBench utilizes StableDiffusion [32] and GPT-4 [1] to generate harmful images and texts based on AdvBench.
@@
 To align with the black-box scenario considered in this paper, we adopt a model transfer strategy, where these white-box methods are trained on one model (LLava-1.5-7b) and then transferred to our target testing models.
-Figure 3: Successful jailbreak instance on Gemini-1.5 using $\mathsf{FC\text{-}Attack}$.
 Defense Settings
@@
 AdaShield-S employs static prompts in the textual prompt to defend against attacks, while AdaShield-A uses Vicunav1.5-13B as a defender to adaptively rewrite defensive prompts [38].
-Figure 4: ASR under different prompts against LVLMs.
+Figure 3: ASR under different prompts against MLLMs.
 Evaluations
 In this section, we explore the performance of $\mathsf{FC\text{-}Attack}$ and conduct ablation study and defense research.
-We conduct jailbreak experiments on LVLMs for $\mathsf{FC\text{-}Attack}$.
-As shown in Figure 3 , it is a successful jailbreak case on Gemini-1.5.
-Figure 5: Different styles of fonts in flowcharts (“1 step”).
-Table 2: Comparison of ASR performance for different models at various description numbers. (Vertical: Top-to-bottom flow; Horizontal: Left-to-right flow; S-shaped: S-shaped path flow.)
+We conduct jailbreak experiments on MLLMs for $\mathsf{FC\text{-}Attack}$.
+As shown in Figure A1 , it is a successful jailbreak case on Gemini-1.5.
+Table 2: ASR comparison across models and attack shapes/sizes.
 Descriptions
@@
 Full 8/8/8/10 8/24/14/30 0/4/0/4 80/76/74/94 76/60/80/92 88/84/88/90 68/60/82/90
+Avg 7/7.5/6/10 7/14.5/13/22 0/2/0/2 78/82/77.5/94 65/54/60/80 81.5/75/78.5/89 65/70.5/68/88
 Table 3: Comparison of ASR (Ensemble) for different font styles and models.
@@
 Performance of $\mathsf{FC\text{-}Attack}$
+Jailbreaking via Images.
 In Table 1, we compare the performance of $\mathsf{FC\text{-}Attack}$ with different baseline methods on both open-source and production models.
-We observe that $\mathsf{FC\text{-}Attack}$ (Ensemble) achieves the highest ASR on both open-source and production models compared to all baselines.
+We observe that $\mathsf{FC\text{-}Attack}$ (Ensemble) achieves the highest ASR on both models compared to all baselines.
 For example, the ASRs are $94\%$, $92\%$, $90\%$, and $90\%$ on Gemini-1.5, Llava-Next, Qwen2-VL, and InternVL-2.5, respectively.
@@
 This difference in performance may indicate that these models struggle to effectively interpret shuffled text and image content.
+Jailbreaking via Videos.
+To conduct attacks from the video modality, we transform each jailbreak image into a 3-second video by setting all frames into the same image.
+Note that we also consider the Procedure Flowcharts, where each part (1 question and 5 steps) has been sequentially filled into a 0.5s video frame, resulting in a 3s video.
+We then evaluate the effectiveness of video jailbreak on three models: Qwen-vl-max, Qwen2.5-Omni and LLaVA-Video.
+The performance is summarized in Table 4.
+Our $\mathsf{FC\text{-}Attack}$ (Ensemble) achieves a stable 88% ASR, whereas HADES peaks at 46% on Qwen-vl-max (dropping to 28% on LLaVA-Video) and Figstep fluctuates between 78% on Qwen-vl-max and 2% on Qwen2.5-Omni, highlighting our method’s consistent performance across models.
+As shown in Figure A2, jailbreaks using harmful text have an extremely low ASR.
+When the same harmful queries and steps are delivered via the video modality, the MLLMs become highly vulnerable, with ASR up to 88%.
+Table 4: Comparison of ASR for different methods and models.
+Method ASR (%)
+Qwen-vl-max Qwen2.5-Omni LLaVA-Video
+HADES 18 40 28
+Figstep 78 2 10
+Ours (Vertical) 72 58 76
+Ours (Ensemble) 88 86 88
+Ours (Procedure) 72 28 82
 Ablation Study
-We then explore the impact of different factors in $\mathsf{FC\text{-}Attack}$ on jailbreak performance, including the different types of user queries, the number of descriptions in flowcharts, and the font styles used in flowcharts.
+We then explore the impact of different factors in $\mathsf{FC\text{-}Attack}$ on jailbreak performance, including the different types of user queries, the number of descriptions, and the font styles used in flowcharts.
 Different Types of User Query.
-We investigate whether the content in flowcharts, when directly input as text, can lead to the jailbroken of LVLMs.
+We investigate whether the content in flowcharts, when directly input as text, can lead to the jailbroken of MLLMs.
 The flowchart content consists of two parts: harmful query from AdvBench and the step descriptions generated by the generator based on this query.
@@
 The ASR is $0\%$ on GPT-4o mini, GPT-4o, Claude-3.5, Qwen2-VL, and InternVL-2.5, and only $2\%$ and $6\%$ on Gemini-1.5 and Llava-Next, respectively.
-This indicates that the textual modality of these LVLMs has relatively robust defenses against such inputs.
+This indicates that the textual modality of these MLLMs has relatively robust defenses against such inputs.
 However, when both the harmful query and the step descriptions are input as text, the ASR increases to $36\%$ on Gemini-1.5, and to $16\%$ and $6\%$ on Llava-Next and InternVL-2.5, respectively, while remaining at $0\%$ on the other models.
 When this information is converted into a flowchart and only a benign textual prompt is provided, the ASR on these models improves significantly.
-This demonstrates that the defenses of LVLMs in the visual modality have noticeable weaknesses compared with the language modality.
-Table 4: Comparison of ASR for different defense methods across various LVLMs.
+This demonstrates that the defenses of MLLMs in the visual modality have noticeable weaknesses compared with the language modality.
+Table 5: Comparison of ASR for different defense methods across various MLLMs.
 Defense ASR (%) (Ensemble)
@@
 A similar trend is observed in Horizontal and S-shaped flowcharts, where ASR reaches $88\%$ and $86\%$ at “5 steps” but decreases to $76\%$ and $74\%$, respectively, at full descriptions.
-This phenomenon may be related to the resolution processing capability of LVLMs.
+This phenomenon may be related to the resolution processing capability of MLLMs.
 When the number of descriptions increases to full, the descriptions may include redundant information, which could negatively impact the model’s performance.
 Font Styles in Flowcharts.
-To investigate whether different font styles in flowcharts affect the effectiveness of jailbreak attacks, we select five fonts from Google Fonts that are relatively difficult for humans to read: Creepster, Fruktur Italic, Pacifico, Shojumaru, and UnifrakturMaguntia (the font style examples are shown in Figure 5).
+To investigate whether different font styles in flowcharts affect the effectiveness of jailbreak attacks, we select five fonts from Google Fonts that are relatively difficult for humans to read: Creepster, Fruktur Italic, Pacifico, Shojumaru, and UnifrakturMaguntia (the font style examples are shown in Figure A3).
 Table 3 shows the results of $\mathsf{FC\text{-}Attack}$ (Ensemble).
@@
 These findings further highlight the need to consider the impact of different font styles when designing defenses.
-Table 5: Performance comparison on the Benign Dataset (MM-vet [35]) under Adashield-S (Ada-S) and Adashield-A (Ada-A) is presented across various LVLMs, reporting scores on six core vision-language capabilities: Recognize (Rec), OCR, Knowledge (Know), Generation (Gen), Spatial (Spat), and Math.
+Table 6: MLLM performance on the Benign MM-Vet dataset [43] under Adashield-S (Ada-S) and Adashield-A (Ada-A), covering six core tasks: Recognize (Rec), OCR, Knowledge (Know), Generation (Gen), Spatial (Spat), and Math.
 Model Benign Dataset Performance (scores)
-[…]
+Defense (rec/ocr/know/gen/spat/math) Total
 GPT4o- mini Vanilla 53.0/68.2/45.7/48.4/60.3/76.5 58.0
@@
 Ada-A 31.5/46.1/19.3/20.9/44.5/41.9 36.7
+Effect of Flowchart Structure.
+To explore the impact of graphical structure elements on the jailbreak effect.
+We conduct experiments with Qwen2-VL using four different flowchart designs: (1) an enhanced FigStep flowchart where each step incorporates step descriptions generated by $\mathsf{FC\text{-}Attack}$; (2) Plain Text structure that only retains text without any graphical elements in the flowchart; (3) Text with Box structure that encapsulates each step in boxes but omits directional arrows; and (4) our complete $\mathsf{FC\text{-}Attack}$ implementation featuring both boxes surrounding step descriptions and arrows indicating the progression between steps.
+Table A1 shows the results of four flowchart image structures.
+We notice that the ASR of the FigStep method is 34%, that of Plain Text is 32%, that of Text with Box is 50%, and that of $\mathsf{FC\text{-}Attack}$ is 88%.
+It is noted that the addition of box elements improves ASR by 18%, while the introduction of directional arrows connecting these boxes further improves it by 38%.
+These findings reveal the contribution of the graphical structural elements of the flowchart to improving the jailbreak effect.
 Defense
-As shown in Table 4, we test four methods to defend against $\mathsf{FC\text{-}Attack}$, where “Original” represents the results of $\mathsf{FC\text{-}Attack}$ (Ensemble) with an average ASR of $58.6\%$.
+We consider four defenses (shown in Table 5), where “Original” represents the results of $\mathsf{FC\text{-}Attack}$ (Ensemble) with an average ASR of $58.6\%$.
 Using Llama-Guard3-V and JailGuard to detect whether the input is harmful reduced the ASR to $51.7\%$ and $51.4\%$, respectively.
-This might be because the information in flowcharts is primarily presented in textual form, while these detection methods are better at identifying information in visual content, making them less effective at detecting textual information.
-AdaShield-S and AdaShield-A reduce the average ASR to $6.9\%$ and $1.7\%$, respectively, by adding defense prompts in the input, showing more effective defense performance.
-However, these two methods also lead to a decline in LVLMs performance on benign datasets.
-We conduct tests on MM-Vet [35] to evaluate the important factor of “over-defensiveness” on benign datasets, which is an evaluation benchmark that contains complex multimodal tasks for LVLMs.
+The limited effectiveness may stem from flowcharts being primarily text-based, whereas the detection methods are more suited to visual content.
+AdaShield-S and AdaShield-A reduce the average ASR to $6.9\%$ and $1.7\%$, showing more effective defense performance.
+However, these two methods also lead to a decline in MLLMs performance on benign datasets.
+We conduct tests on MM-Vet [43] to evaluate the important factor of “over-defensiveness” on benign datasets, which is an evaluation benchmark that contains complex multimodal tasks for MLLMs.
 As shown in Table 6, the model’s utility decreases on benign data when using AdaShield-S and AdaShield-A, indicating a future direction for defense development.
 Conclusion
-In this paper, we propose $\mathsf{FC\text{-}Attack}$, which leverages auto-generated flowcharts to jailbreak LVLMs.
-Experimental results demonstrate that $\mathsf{FC\text{-}Attack}$ achieves higher ASR in both open-source and production LVLMs compared to other jailbreak attack methods.
+In this paper, we propose $\mathsf{FC\text{-}Attack}$, which leverages auto-generated flowcharts to jailbreak MLLMs.
+Experimental results demonstrate that $\mathsf{FC\text{-}Attack}$ achieves higher ASR in both open-source and production MLLMs compared to other jailbreak attacks.
 Additionally, we investigate the factors influencing $\mathsf{FC\text{-}Attack}$, including different types of user queries, the number of steps in flowcharts, and font styles in flowcharts, gaining insights into the aspects that affect ASR.
-Finally, we explore several defense strategies and demonstrate that the AdaShield-A method can effectively mitigate $\mathsf{FC\text{-}Attack}$ but with the cost of utility drop.
+Finally, we explore several defense strategies and demonstrate that the AdaShield-A method can effectively mitigate $\mathsf{FC\text{-}Attack}$, but with the cost of utility drop.
+Limitations
+Our work proposes a novel jailbreak attack on MLLMs via images and videos. However, several limitations remain:
+  • Limited language scope: In this study, we only consider jailbreak attacks conducted in English, as it is the most widely used global language.
+In future work, we plan to explore jailbreak performance in other languages, such as Japanese, Spanish, and Chinese.
+  • Limited model coverage: This work evaluates only 10 representative MLLMs. Future studies can expand this analysis to include more and newer models as they emerge.
+  • Lack of variation in generation parameters: We used a fixed set of generation parameters (e.g., temperature) throughout our experiments.
+We did not investigate how different decoding settings might affect the success of jailbreak attacks.
+We plan to include such analyses in future work.
 Ethical Statement
-This paper presents a method, $\mathsf{FC\text{-}Attack}$, for jailbreaking LVLMs using harmful flowcharts.
-As long as the adversary obtains a harmful flowchart, they can jailbreak LVLMs with minimal resources.
+This paper presents a method, $\mathsf{FC\text{-}Attack}$, for jailbreaking MLLMs using harmful flowcharts.
+As long as the adversary obtains a harmful flowchart, they can jailbreak MLLMs with minimal resources.
 Therefore, it is essential to systematically identify the factors that influence the attack success rate and offer potential defense strategies to model providers.
 Throughout this research, we adhere to ethical guidelines by refraining from publicly distributing harmful flowcharts and harmful responses on the internet before informing service providers of the risks.
-Prior to submitting the paper, we have already sent a warning e-mail to the model providers about the dangers of flowchart-based jailbreak attacks on LVLMs and provided them with the flowcharts generated in our experiments for vulnerability mitigation.
-Appendix A Introduction of LVLMs in this paper
-In this section, we introduce the LVLMs used in this paper.
-  • Llava-Next (January 2024) is an open-source LVLM released by the University of Wisconsin-Madison, which builds upon the Llava-1.5 model [21] with multiple improvements [22].
+Prior to submitting the paper, we have already sent a warning e-mail to the model providers about the dangers of flowchart-based jailbreak attacks on MLLMs and provided them with the flowcharts generated in our experiments for vulnerability mitigation.
+We will release our dataset under the Apache 2.0 License.
+Appendix A Introduction of MLLMs in this paper
+In this section, we introduce the MLLMs used in this paper.
+  • Llava-Next (January 2024) is an open-source MLLM released by the University of Wisconsin-Madison, which builds upon the Llava-1.5 model [25] with multiple improvements [26].
 It enhances capabilities in visual reasoning, optical character recognition, and world knowledge.
@@
 In addition, it adopts multimodal rotary position embedding, effectively integrating positional information across text, images, and videos.
-  • Gemini-1.5 (February 2024) is a production-grade LVLM developed by Google, based on the Mixture-of-Experts architecture [12].
+  • Gemini-1.5 (February 2024) is a production-grade MLLM developed by Google, based on the Mixture-of-Experts architecture [15].
 For Gemini-1.5, larger images will be scaled down to the maximum resolution of $3072\times 3072$, and smaller images will be scaled up to $768\times 768$ pixels.
@@
 The user should submit an image with a long side not larger than 1568 pixels, and the system first scales down the image until it fits the size limit.
-  • GPT-4o and GPT-4o Mini are popular production-grade LVLMs developed by OpenAI [14, 25].
+  • GPT-4o and GPT-4o Mini are popular production-grade MLLMs developed by OpenAI [18, 29].
 GPT-4o Mini is a compact version of GPT-4o, designed for improved cost-efficiency.
 Both models excel in handling complex visual and language understanding tasks.
-  • InternVL-2.5 (June 2024) [10] is an open-source LVLM that ranks first in full-scale open-source multimodal performance.
+  • InternVL-2.5 (June 2024) [13] is an open-source MLLM that ranks first in full-scale open-source multimodal performance.
 In terms of multimodal long-chain reasoning, it achieves a breakthrough of $70\%$ in the expert-level multidisciplinary knowledge reasoning benchmark MMMU [44], and the general capabilities are significantly enhanced.
+  • Qwen-vl-max (January 2024) is the most powerful large-scale visual language model developed by the Alibaba team [5].
+Compared with the enhanced version, the model has made further improvements in visual reasoning and the ability to follow instructions, providing a higher level of visual perception and cognitive understanding.
+It provides the best performance on a wider range of complex tasks, can handle a variety of visual understanding challenges, and demonstrates excellent visual analysis capabilities.
+  • Qwen2.5-Omni (March 2025) is the new flagship end-to-end multimodal model in the Qwen series [41].
+It is designed for comprehensive multimodal perception and can seamlessly handle multiple input forms, including text, images, audio, and video.
+In addition to its powerful multimodal understanding capabilities, Qwen2.5-Omni can also provide real-time streaming responses through text generation and natural speech synthesis.
+  • LLaVA-Video (October 2024) is an extended version of video processing based on the LLaVA-Next model, which inherits the enhanced capabilities of LLaVA-Next in visual reasoning, optical character recognition, and world knowledge, while expanding the processing capabilities of video content [48].
+The model supports high-resolution input (up to 672×672 pixels) and various aspect ratios (336×1344 and 1344×336) to capture more visual details, making it perform well in video understanding and analysis.
 Appendix B Prompts in $\mathsf{FC\text{-}Attack}$
@@
 5. Water regularly without overwatering.
-(a) Flowchart Example-Horizontal: Left-to-right flow
+Appendix D Flowchart Experiment Performance
+Figure A1: Successful jailbreak instance on Gemini-1.5 using $\mathsf{FC\text{-}Attack}$.
+Figure A2: ASR under different prompts against MLLMs via videos.
+Figure A3: Different styles of fonts in flowcharts (“1 step”).
+(a) Flowchart Example-Horizontal: Left-to-right flow.
 (b) Flowchart Example-S-shaped: S-shaped path flow
-(c) Flowchart Example-Vertical: Top-to-bottom flow
-Figure A1: Flowchart Examples: Various flow directions
+(c) Flowchart Example-Vertical: Top-to-bottom flow.
+Figure A4: Flowchart Examples: Various flow directions.
+Table A1: Performances of ASR for different flowchart structures on the Qwen-2-VL.
+Method $\mathsf{FC\text{-}Attack}$ Plain Text Text with Box Figstep
+ASR (%) 88 32 50 34

### v3

@@
 Introduction
-With the advancement of Large Language Models (LLMs), Multimodal Large Language Models (MLLMs) that integrate vision (images and videos) and text, such as GPT-4o [18] and Qwen2.5-VL [6], have demonstrated emergent abilities and achieved impressive performance on downstream tasks related to visual understanding [24, 20].
+With the advancement of Large Language Models (LLMs), Multimodal Large Language Models (MLLMs) that integrate vision (images and videos) and text, such as GPT-4o Hurst et al.
+(2024) and Qwen2.5-VL Bai et al.
+(2025a), have demonstrated emergent abilities and achieved impressive performance on downstream tasks related to visual understanding Liu et al.
+(2024a); Jin et al.
+(2024).
 Figure 1: Comparison of jailbreak effectiveness in various MLLMs using three prompt types.
-Despite being powerful, recent studies [14, 32] have revealed that MLLMs are vulnerable to jailbreak attacks whereby the adversary uses malicious methods to bypass safeguards and gain harmful knowledge.
+Despite being powerful, recent studies Gong et al.
+(2025); Rombach et al.
+(2022) have revealed that MLLMs are vulnerable to jailbreak attacks whereby the adversary uses malicious methods to bypass safeguards and gain harmful knowledge.
 Such vulnerabilities pose remarkable safety risks to the Internet and the physical world.
-For instance, in January 2025, the world witnessed the first case where ChatGPT was used to conduct an explosion [36].
+For instance, in January 2025, the world witnessed the first case where ChatGPT was used to conduct an explosion The Times (2025).
 To better safeguard MLLMs and proactively address their vulnerabilities, model researchers make many efforts in this regard, such as Zhao et al.
-[52] providing a quantitative understanding regarding the adversarial vulnerability of MLLMs.
+(2023b) providing a quantitative understanding regarding the adversarial vulnerability of MLLMs.
 Previous studies often create adversarial datasets tailored to specific models, which tend to perform poorly on other models.
-Currently, jailbreak attacks against MLLMs can be broadly categorized into two main types: optimization-based attacks [8, 23] and prompt-based attacks [14, 39].
+Currently, jailbreak attacks against MLLMs can be broadly categorized into two main types: optimization-based attacks Bailey et al.
+(2024); Li et al.
+(2024b) and prompt-based attacks Gong et al.
+(2025); Wang et al.
+(2025).
 Optimization-based attacks use white-box gradient methods to craft adversarial perturbations on visual prompt aligned with harmful text.
@@
 Concretely, $\mathsf{FC\text{-}Attack}$ converts harmful queries into harmful flowcharts (images and videos) as visual prompts, allowing users to input benign textual prompts to bypass the model’s safeguards.
-Specifically, $\mathsf{FC\text{-}Attack}$ consists of two stages: (1) Step-Description Generator Building: In this stage, the step description dataset is synthesized using GPT-4o, and fine-tune a pre-trained LLM to obtain a step-description generator.
+Specifically, $\mathsf{FC\text{-}Attack}$ consists of two stages: (1) Step-Description Generator Building: In this stage, the step description dataset is synthesized using GPT-4o, and a pre-trained LLM is fine-tuned to obtain a step-description generator.
 (2) Jailbreak Deployment: This stage uses the generator to produce steps corresponding to the harmful query and generates three types of harmful flowcharts (vertical, horizontal, and S-shaped) as visual prompts.
@@
 Note that the harmful flowcharts are generated automatically without hand-crafted effort.
-Our evaluation on the Advbench dataset shows that $\mathsf{FC\text{-}Attack}$ outperforms previous attacks and achieves an attack success rate (ASR) of over $90\%$ on multiple open-source models, including Llava-Next, Qwen2-VL, and InternVL-2.5, and reaches $94\%$ on the production model Gemini-1.5.
+Our evaluation on the Advbench dataset Zou et al.
+(2023) shows that $\mathsf{FC\text{-}Attack}$ outperforms previous attacks and achieves an attack success rate (ASR) of over $90\%$ on multiple open-source models, including Llava-Next, Qwen2-VL, and InternVL-2.5, and reaches $94\%$ on the production model Gemini-1.5.
 Although the ASR is lower on GPT-4o mini, GPT-4o, and Claude-3.5, we how later that it can be improved in certain ways.
-To further investigate the impact of different elements in flowcharts on the jailbreak effectiveness of MLLMs, we conduct several ablation experiments, including different types of user queries (as shown in Figure 1), numbers of descriptions, and font styles in flowcharts.
+To further investigate the impact of different elements in flowcharts on the jailbreak effectiveness of MLLMs, we conduct several ablation experiments, including different types of user queries (as shown in Figure˜1), numbers of descriptions, and font styles in flowcharts.
 These experiments show that MLLMs exhibit higher safety in the text modality but weaker in the visual modality.
@@
 For instance, when the font style is changed from “Times New Roman” to “Pacifico”, the ASR increases from $4\%$ to $28\%$ on the model with the lowest ASR (Claude-3.5) under the original style.
-To mitigate the attack, we consider several popular defense approaches, including Llama-Guard-3-11B-Vision [28], JailGuard [47], AdaShield-S [38], and AdaShield-A [38].
+To mitigate the attack, we consider several popular defense approaches, including Llama-Guard-3-11B-Vision Meta LLaMA (2025), JailGuard Zhang et al.
+(2024b), AdaShield-S Wang et al.
+(2024b), and AdaShield-A Wang et al.
+(2024b).
 Among them, AdaShield-A demonstrates the best defense performance by reducing the average ASR from $58.6\%$ to $1.7\%$.
@@
 Multimodal Large Language Models
-In recent years, with the increase in model parameters and training data, LLMs have demonstrated powerful language generation and understanding capabilities [51, 10], which have driven the emergence of MLLMs [46] (also known as Large Vision Language Models, LVLMs).
-MLLMs combine visual understanding with language comprehension, showing promising capabilities in visual downstream tasks, including Visual Question Answering (VQA) [3, 21, 33], image captioning [17, 22], and visual commonsense reasoning [45, 35].
+In recent years, with the increase in model parameters and training data, LLMs have demonstrated powerful language generation and understanding capabilities Zhao et al.
+(2023a); Chang et al.
+(2024), which have driven the emergence of MLLMs Zhang et al.
+(2024a) (also known as Large Vision Language Models, LVLMs).
+MLLMs combine visual understanding with language comprehension, showing promising capabilities not only in canonical visual tasks such as Visual Question Answering (VQA) Antol et al.
+(2015); Khan et al.
+(2023); Shao et al.
+(2023), image captioning Hu et al.
+(2021); Li et al.
+(2024a), and visual commonsense reasoning Zellers et al.
+(2019); Tanaka et al.
+(2021), but also in emerging applications including edited image detection Sun et al.
+(2025b) and real-time assistance for individuals with visual impairments Zhang et al.
+(2025).
 Notably, some MLLMs are capable of processing both image and video inputs, enabling broader applications across multimodal scenarios.
-In this paper, we consider both popular open-source and widely used production MLLMs.
+In this paper, we consider both popular open-source and production MLLMs.
 These MLLMs are the most widely used, and all of them have been aligned to ensure safety.
-Detailed information are introduced in Appendix A.
+Detailed information is introduced in Appendix˜A.
 Figure 2: Overview of the $\mathsf{FC\text{-}Attack}$ framework with two stages.
 Jailbreak Attacks on MLLMs
-Similar to LLMs, which have been shown to be vulnerable to jailbreak attacks [42], MLLMs also remain susceptible despite safety alignment.
+Similar to LLMs, which have been shown to be vulnerable to jailbreak attacks Yi et al.
+(2024), MLLMs also remain susceptible despite safety alignment.
 Current attacks can be categorized into two types: optimization-based and prompt-based attacks.
 Most existing optimization-based attacks rely on backpropagating the gradient of the target to generate harmful outputs.
-These methods typically require white-box access to the model, where they obtain the output logits of MLLMs and then compute the loss with the target response to create adversarial perturbations into the visual prompts or textual prompts [4, 34, 31] (e.g., the target can be “Sure!
+These methods typically require white-box access to the model, where they obtain the output logits of MLLMs and then compute the loss with the target response to create adversarial perturbations into the visual prompts or textual prompts Bagdasaryan et al.
+(2023); Shayegani et al.
+(2024); Qi et al.
+(2024) (e.g., the target can be “Sure!
 I’m ready to answer your question.”).
 Carlini et al.
-[9] are the first to propose optimizing input images by using fixed toxic outputs as targets, thereby forcing the model to produce harmful outputs.
+(2023) are the first to propose optimizing input images by using fixed toxic outputs as targets, thereby forcing the model to produce harmful outputs.
 Building on this, Bailey et al.
-[8] introduce the Behaviour Matching Algorithm, which trains adversarial images to make MLLMs output behavior that matches a target in specific contextual inputs.
+(2024) introduce the Behaviour Matching Algorithm, which trains adversarial images to make MLLMs output behavior that matches a target in specific contextual inputs.
 This process requires the model’s output logits to align closely with those of the target behavior.
@@
 Li et al.
-[23] take this further by replacing harmful keywords in the original textual inputs with objects or actions in the image, allowing harmful information to be conveyed through images to achieve jailbreaking.
-Unlike previous work, these images are generated using diffusion models and are iteratively optimized with models like GPT-4.
+(2024b) take this further by replacing harmful keywords in the original textual inputs with objects or actions in the image, allowing harmful information to be conveyed through images to achieve jailbreaking.
+Unlike previous work, these images are generated using diffusion models and are iteratively optimized with models like GPT-4 Achiam et al.
+(2023).
 This approach enhances the harmfulness of the images, enabling more effective attacks.
@@
 Gong et al.
-[14] discovers that introducing visual modules may cause the original security mechanisms of LLMs to fail in covering newly added visual content, resulting in potential security vulnerabilities.
+(2025) discovers that introducing visual modules may cause the original security mechanisms of LLMs to fail in covering newly added visual content, resulting in potential security vulnerabilities.
 To address this, they propose the FigStep attack, which converts harmful textual instructions into text embedded in images and uses a neutral textual prompt to guide the model into generating harmful content.
@@
 Wang et al.
-[39] identifies a phenomenon named Shuffle Inconsistency, which highlights the tension between “understanding capabilities” and “safety mechanisms” of LLMs.
+(2025) identifies a phenomenon named Shuffle Inconsistency, which highlights the tension between “understanding capabilities” and “safety mechanisms” of LLMs.
 Specifically, even if harmful instructions in text or images are rearranged, MLLMs can still correctly interpret their meaning.
@@
 Adversary’s Goal.
-The adversary’s goal is to exploit attacks to bypass the protective mechanisms of MLLMs and access content prohibited by safety policies, e.g., OpenAI’s usage policy [30].
+The adversary’s goal is to exploit attacks to bypass the protective mechanisms of MLLMs and access content prohibited by safety policies, e.g., OpenAI’s usage policy OpenAI (2025).
 This goal takes real-world scenarios into account, where adversaries manipulate the capabilities of MLLMs to easily acquire harmful knowledge and thereby commit criminal acts with minimal learning effort.
@@
 Our Method
-In this section, we introduce the framework of $\mathsf{FC\text{-}Attack}$ (as shown in Figure 2), which consists of two stages: Step-Description Generator Building and Jailbreak Deployment.
+In this section, we introduce the framework of $\mathsf{FC\text{-}Attack}$ (as shown in Figure˜2), which consists of two stages: Step-Description Generator Building and Jailbreak Deployment.
 Step-Description Generator Building
@@
 After obtaining the Step-Description Generator $\mathcal{G}$, a harmful query $q_{h}$ is input to generate the corresponding step-by-step description.
-This description is then processed by a transformation function $\mathcal{F}$ to generate the flowchart (using Graphviz [16]).
-Together with a benign textual prompt $p_{b}$ (more details are in Appendix B), the flowchart will be fed into the aligned MLLM $\mathcal{A}$ to produce the harmful output $o_{h}$, as shown below:
+This description is then processed by a transformation function $\mathcal{F}$ to generate the flowchart (using Graphviz Graphviz Team (2025)).
+Together with a benign textual prompt $p_{b}$ (more details are in Appendix˜B), the flowchart will be fed into the aligned MLLM $\mathcal{A}$ to produce the harmful output $o_{h}$, as shown below:
 $o_{h}=\mathcal{A}(\mathcal{F}(\mathcal{G}(q_{h})),p_{b})\leftarrow\text{FC-attack}(q_{h}).$ (4)
@@
 Target Model.
-We test $\mathsf{FC\text{-}Attack}$ on seven popular MLLMs, including the open-source models Llava-Next ( llama3-llava-next-8b) [26], Qwen2-VL (Qwen2-VL-7B-Instruct) [37], and InternVL-2.5 (InternVL-2.5-8B) [12] as well as the production models GPT-4o mini (gpt-4o-mini-2024-07-18) [29], GPT-4o (gpt-4o-2024-08-06) [18], Claude (claude-3-5-sonnet-20240620) [2], and Gemini (gemini-1.5-flash) [15].
-Moreover, we also test $\mathsf{FC\text{-}Attack}$ on LLMs via video, including Qwen-vl-max (Qwen-vl-max-latest) [7],Qwen2.5-Omni [40] and LLaVA-Video-7B-Qwen2 [49].
+We test $\mathsf{FC\text{-}Attack}$ on seven popular MLLMs, including the open-source models Llava-Next (llama3-llava-next-8b) Liu et al.
+(2024c), Qwen2-VL (Qwen2-VL-7B-Instruct) Wang et al.
+(2024a), and InternVL-2.5 (InternVL-2.5-8B) Chen et al.
+(2024a) as well as the production models GPT-4o mini (gpt-4o-mini-2024-07-18) OpenAI (2024), GPT-4o (gpt-4o-2024-08-06) Hurst et al.
+(2024), Claude (claude-3-5-sonnet-20240620) Anthropic (2024), and Gemini (gemini-1.5-flash) Google (2024).
+Moreover, we also test $\mathsf{FC\text{-}Attack}$ on MLLMs via video, including Qwen-VL-Max (Qwen-VL-Max-latest) Bai et al.
+(2025b), Qwen2.5-Omni Xu et al.
+(2025) and LLaVA-Video-7B-Qwen2 Zhang et al.
+(2024d).
 Dataset.
 Following Chao et al.
-[11], we utilize the deduplicated version of AdvBench [54], which includes 50 representative harmful queries.
+(2025b), we utilize the deduplicated version of Advbench Zou et al.
+(2023), which includes 50 representative harmful queries.
 Based on Advbench, we use $\mathsf{FC\text{-}Attack}$ to generate $3$ types of flowcharts for each harmful query, which includes $150$ jailbreak flowcharts in total.
-To assess whether defense methods have the critical issue of “over-defensiveness” when applied to benign datasets, we utilize a popular evaluation benchmark, MM-Vet [43].
+To assess whether defense methods have the critical issue of “over-defensiveness” when applied to benign datasets, we utilize a popular evaluation benchmark, MM-Vet Yu et al.
+(2024).
 Table 1: Comparison of ASR performance across different methods and MLLMs.
@@
 GPT-4o mini GPT-4o Claude-3.5 Gemini-1.5 Llava-Next Qwen2-VL InternVL-2.5
-HADES $4$ $16$ $2$ $2$ $20$ $10$ $8$
-SI-Attack $36$ $14$ 00 $69$ $24$ $42$ $40$
-MM-SafetyBench 00 00 00 $50$ $50$ $54$ $16$
-VA-Jailbreak $6$ $18$ $2$ $2$ $40$ $22$ $16$
-FigStep 00 $2$ 00 $30$ $62$ $36$ 00
-Ours (Vertical) $8$ $8$ 00 $76$ $76$ $84$ $68$
+HADES 44 $16$ 22 22 2020 $10$ 88
+SI-Attack $36$ $14$ 0 $69$ $24$ $42$ $40$
+$M$-SafetyBench 0 0 0 $50$ $50$ $54$ $16$
+VA-Jailbreak 66 $18$ 22 22 $40$ $22$ $16$
+FigStep 0 22 0 $30$ $62$ $36$ 0
+Ours (Vertical) 88 88 0 $76$ $76$ $84$ $68$
 Ours (Ensemble) $\mathbf{10}$ $\mathbf{30}$ $\mathbf{4}$ $\mathbf{94}$ $\mathbf{92}$ $\mathbf{90}$ $\mathbf{90}$
 Evaluation Metric. In the experiments, we use the ASR to evaluate the performance of our attack, which can be defined as follows:
-ASR=# Queries Successfully Jailbroken# Original Harmful $\text{ASR}=\frac{\text{\# Queries Successfully Jailbroken}}{\text{\# Original Harmful Queries}}.$ (5)
-Following the judge prompt [11], we employ GPT-4o to serve as the evaluator.
+$\text{ASR}=\frac{\text{\# Queries Successfully Jailbroken}}{\text{\# Original Harmful Queries}}.$ (5)
+Following the judge prompt Chao et al. (2025b), we employ GPT-4o to serve as the evaluator.
 FC-Attack Deployment.
-Referring to Section 4, $\mathsf{FC\text{-}Attack}$ consists of two stages.
+Referring to Section˜4, $\mathsf{FC\text{-}Attack}$ consists of two stages.
 For the Step-Description Generator Building, we first use GPT-4o to randomly generate several daily topics and $3$ few-shot examples, which are then combined into a prompt and fed into GPT-4o to construct the dataset $\mathcal{D}_{t}$.
 In our experiments, the number of descriptions in the flowchart is limited to a maximum of $10$ steps, as too many descriptions can result in excessive length in one direction of the image.
-The dataset contains $5,000$ pairs of queries and step descriptions for daily activities, with the temperature set to 1 (more details are provided in Appendix C).
-We then select Mistral-7B-Instruct-v0.1 [19] as the pre-trained LLM and fine-tune it on $\mathcal{D}_{t}$ using LoRA.
-The fine-tuning parameters include a rank of $16$, a LoRA alpha value of $64$, $2$ epochs, a batch size of $8$, a learning rate of $51e-$ 5, and a weight decay of $51e-$ 5.
+The dataset contains $5,000$ pairs of queries and step descriptions for daily activities, with the temperature set to 1 (more details are provided in Appendix˜C).
+We then select Mistral-7B-Instruct-v0.1 Jiang et al.
+(2023) as the pre-trained LLM and fine-tune it on $\mathcal{D}_{t}$ using LoRA.
+The fine-tuning parameters include a rank of $16$, a LoRA alpha value of $64$, 22 epochs, a batch size of $8$, a learning rate of 1e51e-5, and a weight decay of 1e51e-5.
 For the jailbreak deployment stage, we set the temperature to $0.3$ for all MLLMs for a fair comparison.
 Baselines.
-To validate the effectiveness of $\mathsf{FC\text{-}Attack}$, we adopt five jailbreak attacks as baselines, which are categorized into black-box attacks (MM-SafetyBench [27], SI-Attack [50], and FigStep [14]) and white-box attacks (HADES [23], VA-Jailbreak [31]).
-For black-box attacks, MM-SafetyBench utilizes StableDiffusion [32] and GPT-4 [1] to generate harmful images and texts based on AdvBench.
-The input harmful images and texts used in SI-Attack are from the outputs of MM-SafetyBench, while FigStep is set up using their default settings [14].
+To validate the effectiveness of $\mathsf{FC\text{-}Attack}$, we adopt five jailbreak attacks as baselines, which are categorized into black-box attacks (MM-SafetyBench Liu et al.
+(2024d), SI-Attack Zhao et al.
+(2025), and FigStep Gong et al.
+(2025)) and white-box attacks (HADES Li et al.
+(2024b), VA-Jailbreak Qi et al.
+(2024)).
+For black-box attacks, MM-SafetyBench utilizes StableDiffusion Rombach et al.
+(2022) and GPT-4 Achiam et al.
+(2023) to generate harmful images and texts based on Advbench.
+The input harmful images and texts used in SI-Attack are from the outputs of MM-SafetyBench, while FigStep is set up using their default settings Gong et al.
+(2025).
 For white-box attacks, all input data, including images and texts, is obtained from MM-SafetyBench’s outputs, with the attack step size uniformly set to 1/255.
-HADES employs LLaVa-1.5-7b [25] as the attack model, running 3,000 optimization iterations with a batch size of 2.
-For VA-Jailbreak, LLaVa-1.5-7b [25] is used as the attack model, setting the epsilon of the attack budget to 32/255, with 5,000 optimization iterations and a batch size of 8.
+HADES employs LLaVa-1.5-7b Liu et al.
+(2024b) as the attack model, running 3,000 optimization iterations with a batch size of 2.
+For VA-Jailbreak, LLaVa-1.5-7b Liu et al.
+(2024b) is used as the attack model, setting the epsilon of the attack budget to 32/255, with 5,000 optimization iterations and a batch size of 8.
 To align with the black-box scenario considered in this paper, we adopt a model transfer strategy, where these white-box methods are trained on one model (LLava-1.5-7b) and then transferred to our target testing models.
 Defense Settings
-To mitigate the attacks, we explore several possible defense methods including Llama-Guard3-V, JailGuard, and AdaShield.
-Llama-Guard3-V (Llama-Guard-3-11B-Vision) [28] determines whether the input is safe by feeding both the image and text into the model.
-JailGuard [47] generates input variants and evaluates them using MiniGPT-4 [53], identifying harmful content by comparing differences in the responses.
-AdaShield-S employs static prompts in the textual prompt to defend against attacks, while AdaShield-A uses Vicunav1.5-13B as a defender to adaptively rewrite defensive prompts [38].
-Figure 3: ASR under different prompts against MLLMs.
+To mitigate the attacks, we explore several possible defense methods including Llama-Guard3-V, JailGuard, ECSO and AdaShield.
+Llama-Guard3-V (Llama-Guard-3-11B-Vision) Meta LLaMA (2025) determines whether the input is safe by feeding both the image and text into the model.
+JailGuard Zhang et al.
+(2024b) generates input variants and evaluates them using MiniGPT-4 Zhu et al.
+(2024), identifying harmful content by comparing differences in the responses.
+ECSO Chao et al.
+(2025a) is less intrusive but mainains a high safety rate simultaneously.
+AdaShield-S employs static prompts in the textual prompt to defend against attacks, while AdaShield-A uses Vicunav1.5-13B as a defender to adaptively rewrite defensive prompts Wang et al.
+(2024b).
 Evaluations
@@
 We conduct jailbreak experiments on MLLMs for $\mathsf{FC\text{-}Attack}$.
-As shown in Figure A1 , it is a successful jailbreak case on Gemini-1.5.
+As shown in Figure˜3 , it is a successful jailbreak case on Gemini-1.5.
+Figure 3: Successful jailbreak instance on Gemini-1.5 using $\mathsf{FC\text{-}Attack}$.
 Table 2: ASR comparison across models and attack shapes/sizes.
@@
 Original 10 30 4 94 92 90 90
-Creepster 14$\uparrow$ 24$\downarrow$ 8$\uparrow$ 94 90$\downarrow$ 90 90
-Fruktur 18$\uparrow$ 28$\downarrow$ 18$\uparrow$ 98$\uparrow$ 86$\downarrow$ 90 88$\downarrow$
-Pacifico 14$\uparrow$ 30 28$\uparrow$ 90$\downarrow$ 90$\downarrow$ 90 96$\uparrow$
-Shojumaru 20$\uparrow$ 30 12$\uparrow$ 90$\downarrow$ 94$\uparrow$ 90 88$\downarrow$
-UnifrakturMaguntia 12$\uparrow$ 24$\downarrow$ 26$\uparrow$ 90$\downarrow$ 90$\downarrow$ 90 92$\uparrow$
+[…]
+[…]
+[…]
+[…]
+[…]
 Performance of $\mathsf{FC\text{-}Attack}$
 Jailbreaking via Images.
-In Table 1, we compare the performance of $\mathsf{FC\text{-}Attack}$ with different baseline methods on both open-source and production models.
+In Table˜1, we compare the performance of $\mathsf{FC\text{-}Attack}$ with different baseline methods on both open-source and production models.
 We observe that $\mathsf{FC\text{-}Attack}$ (Ensemble) achieves the highest ASR on both models compared to all baselines.
@@
 We then evaluate the effectiveness of video jailbreak on three models: Qwen-VL-Max, Qwen2.5-Omni and LLaVA-Video.
-The performance is summarized in Table 4.
+The performance is summarized in Table˜4.
 Our $\mathsf{FC\text{-}Attack}$ (Ensemble) achieves a stable 88% ASR, whereas HADES peaks at 46% on Qwen-VL-Max (dropping to 28% on LLaVA-Video) and Figstep fluctuates between 78% on Qwen-VL-Max and 2% on Qwen2.5-Omni, highlighting our method’s consistent performance across models.
-As shown in Figure A2, jailbreaks using harmful text have an extremely low ASR.
+As shown in Figure˜A2, attacks using harmful text have an extremely low ASR.
 When the same harmful queries and steps are delivered via the video modality, the MLLMs become highly vulnerable, with ASR up to 88%.
@@
 The flowchart content consists of two parts: harmful query from Advbench and the step descriptions generated by the generator based on this query.
-As shown in Figure 3 when using only the harmful query (text) as input, we observe very low ASR.
+Figure 4: ASR under different prompts against MLLMs.
+As shown in Figure˜4 when using only the harmful query (text) as input, we observe very low ASR.
 The ASR is $0\%$ on GPT-4o mini, GPT-4o, Claude-3.5, Qwen2-VL, and InternVL-2.5, and only $2\%$ and $6\%$ on Gemini-1.5 and Llava-Next, respectively.
@@
 JailGuard 8 24 2 86 80 82 78 51.4
+ECSO 2 0 0 42 44 30 42 22.6
 AdaShield-S 0 0 0 12 22 10 4 6.9
@@
 Numbers of Steps in Flowcharts.
-As described in Section 4, flowcharts of $\mathsf{FC\text{-}Attack}$ are generated from step descriptions.
+As described in Section˜4, flowcharts of $\mathsf{FC\text{-}Attack}$ are generated from step descriptions.
 In this section, we aim to explore the impact of the number of steps in flowcharts on jailbreak effectiveness.
 Therefore, we reduce the number of steps to 1, 3, and 5, respectively.
-Table 2 presents the ASR results for four types of flowcharts (Vertical, Horizontal, S-shaped, and Ensemble) with varying numbers of steps.
+Table˜2 presents the ASR results for four types of flowcharts (Vertical, Horizontal, S-shaped, and Ensemble) with varying numbers of steps.
 We find that, even with only one step in the description, flowcharts achieve relatively high ASR.
@@
 Font Styles in Flowcharts.
-To investigate whether different font styles in flowcharts affect the effectiveness of jailbreak attacks, we select five fonts from Google Fonts that are relatively difficult for humans to read: Creepster, Fruktur Italic, Pacifico, Shojumaru, and UnifrakturMaguntia (the font style examples are shown in Figure A3).
-Table 3 shows the results of $\mathsf{FC\text{-}Attack}$ (Ensemble).
+To investigate whether different font styles in flowcharts affect the effectiveness of jailbreak attacks, we select five fonts from Google Fonts that are relatively difficult for humans to read: Creepster, Fruktur Italic, Pacifico, Shojumaru, and UnifrakturMaguntia (the font style examples are shown in Figure˜5).
+Table˜3 shows the results of $\mathsf{FC\text{-}Attack}$ (Ensemble).
 We observe that different font styles can significantly impact the ASR.
@@
 These findings further highlight the need to consider the impact of different font styles when designing defenses.
-Table 6: MLLM performance on the Benign MM-Vet dataset [43] under Adashield-S (Ada-S) and Adashield-A (Ada-A), covering six core tasks: Recognize (Rec), OCR, Knowledge (Know), Generation (Gen), Spatial (Spat), and Math.
+Figure 5: Different styles of fonts in flowcharts (“1 step”).
+Table 6: MLLM performance on the benign $M$-Vet dataset Yu et al.
+(2024) under Adashield-S (Ada-S) and Adashield-A (Ada-A), covering six core tasks: Recognize (Rec), OCR, Knowledge (Know), Generation (Gen), Spatial (Spat), and Math.
 Model Benign Dataset Performance (scores)
 Defense (rec/ocr/know/gen/spat/math) Total
-GPT4o- mini Vanilla 53.0/68.2/45.7/48.4/60.3/76.5 58.0
+GPT4o-
+mini
+Vanilla 53.0/68.2/45.7/48.4/60.3/76.5 58.0
 Ada-S 35.1/66.7/30.4/34.1/55.7/76.5 45.1
@@
 Ada-A 59.5/74.3/56.1/58.9/67.9/83.1 64.6
-Claude- 3.5 Vanilla 61.1/72.8/51.8/52.0/70.7/80.0 64.8
+Claude-
+3.5
+Vanilla 61.1/72.8/51.8/52.0/70.7/80.0 64.8
 Ada-S 60.1/69.7/50.1/51.5/66.9/75.4 62.8
 Ada-A 59.5/70.6/52.5/51.7/67.5/74.2 63.2
-Gemini- 1.5 Vanilla 59.9/73.7/50.8/50.9/69.5/85.4 64.2
+Gemini-
+1.5
+Vanilla 59.9/73.7/50.8/50.9/69.5/85.4 64.2
 Ada-S 53.8/69.6/43.7/43.6/66.8/75.4 58.2
 Ada-A 54.8/72.6/44.2/44.0/69.3/81.2 59.9
-Llava- Next Vanilla 38.0/39.0/25.8/24.8/40.1/21.2 38.8
+Llava-
+Next
+Vanilla 38.0/39.0/25.8/24.8/40.1/21.2 38.8
 Ada-S 33.7/42.0/26.7/25.1/43.7/36.2 37.0
 Ada-A 36.5/37.7/24.8/24.3/37.6/18.8 36.7
-Qwen2- VL Vanilla 51.9/62.4/44.5/41.6/55.5/60.4 55.0
+Qwen2-
+VL
+Vanilla 51.9/62.4/44.5/41.6/55.5/60.4 55.0
 Ada-S 39.3/55.0/31.1/29.1/50.5/46.2 44.9
 Ada-A 44.5/57.5/34.2/33.2/55.7/58.8 49.8
-InternVL- 2.5 Vanilla 52.0/55.4/42.6/40.1/55.6/45.4 53.1
+InternVL-
+2.5
+Vanilla 52.0/55.4/42.6/40.1/55.6/45.4 53.1
 Ada-S 27.2/43.2/16.4/20.2/40.3/45.8 31.9
@@
 We conduct experiments with Qwen2-VL using four different flowchart designs: (1) an enhanced FigStep flowchart where each step incorporates step descriptions generated by $\mathsf{FC\text{-}Attack}$; (2) Plain Text structure that only retains text without any graphical elements in the flowchart; (3) Text with Box structure that encapsulates each step in boxes but omits directional arrows; and (4) our complete $\mathsf{FC\text{-}Attack}$ implementation featuring both boxes surrounding step descriptions and arrows indicating the progression between steps.
-Table A1 shows the results of four flowchart image structures.
-We notice that the ASR of the FigStep method is 34%, that of Plain Text is 32%, that of Text with Box is 50%, and that of $\mathsf{FC\text{-}Attack}$ is 88%.
+Table˜A1 shows the results of four flowchart image structures.
+We notice that the ASR of the FigStep method is 34%, that of Plain Text is 32%, that of Text with Box is 50%, and that of $\mathsf{FC\text{-}Attack}$ is 90%.
 It is noted that the addition of box elements improves ASR by 18%, while the introduction of directional arrows connecting these boxes further improves it by 38%.
 These findings reveal the contribution of the graphical structural elements of the flowchart to improving the jailbreak effect.
-Defense
-We consider four defenses (shown in Table 5), where “Original” represents the results of $\mathsf{FC\text{-}Attack}$ (Ensemble) with an average ASR of $58.6\%$.
+Effect of Different Formats.
+To investigate the impact of different formats on ASR, we add an experiment on three different formats (Code, Figstep-Style presentations, and Table).
+All methods deliver the same harmful steps but present them in different formats.
+We conduct the experiment on the Qwen2-VL model.
+(1) Code: Python code form; (2) Figstep-Style presentations: Fill in the same full steps into the Figstep-Style format; (3) Table: Fill in the same full steps into the table format.
+From table˜A2, we can observe that Figstep-Style performs only 12% ASR, followed by the Table format at 26% ASR.
+The Code format shows moderate effectiveness with 52% ASR.
+In contrast, our $\mathsf{FC\text{-}Attack}$ achieves the highest ASR at 90%, which indicates flowchart is the most effective format.
+Defenses
+We consider four defenses (shown in Table˜5), where “Original” represents the results of $\mathsf{FC\text{-}Attack}$ (Ensemble) with an average ASR of $58.6\%$.
 Using Llama-Guard3-V and JailGuard to detect whether the input is harmful reduced the ASR to $51.7\%$ and $51.4\%$, respectively.
@@
 However, these two methods also lead to a decline in MLLMs performance on benign datasets.
-We conduct tests on MM-Vet [43] to evaluate the important factor of “over-defensiveness” on benign datasets, which is an evaluation benchmark that contains complex multimodal tasks for MLLMs.
-As shown in Table 6, the model’s utility decreases on benign data when using AdaShield-S and AdaShield-A, indicating a future direction for defense development.
+We conduct tests on MM-Vet Yu et al.
+(2024) to evaluate the important factor of “over-defensiveness” on benign datasets, which is an evaluation benchmark that contains complex multimodal tasks for MLLMs.
+As shown in Table˜6, the model’s utility decreases on benign data when using AdaShield-S and AdaShield-A, indicating a future direction for defense development.
 Conclusion
@@
 We will release our dataset under the Apache 2.0 License.
+Acknowledgments
+We would like to thank all anonymous reviewers, Area Chair, and Program Chair for their insightful comments and constructive suggestions.
 Appendix A Introduction of MLLMs in this paper
 In this section, we introduce the MLLMs used in this paper.
-  • Llava-Next (January 2024) is an open-source MLLM released by the University of Wisconsin-Madison, which builds upon the Llava-1.5 model [25] with multiple improvements [26].
+  • Llava-Next (January 2024) is an open-source MLLM released by the University of Wisconsin-Madison, which builds upon the Llava-1.5 model Liu et al.
+[2024b] with multiple improvements Liu et al.
+[2024c].
 It enhances capabilities in visual reasoning, optical character recognition, and world knowledge.
 Besides, Llava-Next increases the input image resolution to a maximum of $672\times 672$ pixels and supports various aspect ratios to capture more visual details ($336\times 1344$ and $1344\times 336$).
-  • Qwen2-VL (September 2024) is an open-source model released by Alibaba team [37].
+  • Qwen2-VL (September 2024) is an open-source model released by the Alibaba team Wang et al.
+[2024a].
 It employs naive dynamic resolution to handle images of different resolutions.
 In addition, it adopts multimodal rotary position embedding, effectively integrating positional information across text, images, and videos.
-  • Gemini-1.5 (February 2024) is a production-grade MLLM developed by Google, based on the Mixture-of-Experts architecture [15].
+  • Gemini-1.5 (February 2024) is a production-grade MLLM developed by Google, based on the Mixture-of-Experts architecture Google [2024].
 For Gemini-1.5, larger images will be scaled down to the maximum resolution of $3072\times 3072$, and smaller images will be scaled up to $768\times 768$ pixels.
 Reducing the image size will not improve the performance of higher-resolution images.
-  • Claude-3.5-Sonnet (June 2024) is a production multimodal AI assistant developed by Anthropic [2].
+  • Claude-3.5-Sonnet (June 2024) is a production multimodal AI assistant developed by Anthropic Anthropic [2024].
 The user should submit an image with a long side not larger than 1568 pixels, and the system first scales down the image until it fits the size limit.
-  • GPT-4o and GPT-4o Mini are popular production-grade MLLMs developed by OpenAI [18, 29].
+  • GPT-4o and GPT-4o Mini are popular production-grade MLLMs developed by OpenAI Hurst et al.
+[2024], OpenAI [2024].
 GPT-4o Mini is a compact version of GPT-4o, designed for improved cost-efficiency.
 Both models excel in handling complex visual and language understanding tasks.
-  • InternVL-2.5 (June 2024) [13] is an open-source MLLM that ranks first in full-scale open-source multimodal performance.
-In terms of multimodal long-chain reasoning, it achieves a breakthrough of $70\%$ in the expert-level multidisciplinary knowledge reasoning benchmark MMMU [44], and the general capabilities are significantly enhanced.
-  • Qwen-vl-max (January 2024) is the most powerful large-scale visual language model developed by the Alibaba team [5].
+  • InternVL-2.5 (June 2024) Chen et al.
+[2024b] is an open-source MLLM that ranks first in full-scale open-source multimodal performance.
+In terms of multimodal long-chain reasoning, it achieves a breakthrough of $70\%$ in the expert-level multidisciplinary knowledge reasoning benchmark MMMU Yue et al.
+[2024], and the general capabilities are significantly enhanced.
+  • Qwen-VL-Max (January 2024) is the most powerful large-scale visual language model developed by the Alibaba team Bai et al.
+[2023].
 Compared with the enhanced version, the model has made further improvements in visual reasoning and the ability to follow instructions, providing a higher level of visual perception and cognitive understanding.
 It provides the best performance on a wider range of complex tasks, can handle a variety of visual understanding challenges, and demonstrates excellent visual analysis capabilities.
-  • Qwen2.5-Omni (March 2025) is the new flagship end-to-end multimodal model in the Qwen series [41].
+  • Qwen2.5-Omni (March 2025) is the new flagship end-to-end multimodal model in the Qwen series Yang et al.
+[2024].
 It is designed for comprehensive multimodal perception and can seamlessly handle multiple input forms, including text, images, audio, and video.
 In addition to its powerful multimodal understanding capabilities, Qwen2.5-Omni can also provide real-time streaming responses through text generation and natural speech synthesis.
-  • LLaVA-Video (October 2024) is an extended version of video processing based on the LLaVA-Next model, which inherits the enhanced capabilities of LLaVA-Next in visual reasoning, optical character recognition, and world knowledge, while expanding the processing capabilities of video content [48].
+  • LLaVA-Video (October 2024) is an extended version of video processing based on the LLaVA-Next model, which inherits the enhanced capabilities of LLaVA-Next in visual reasoning, optical character recognition, and world knowledge, while expanding the processing capabilities of video content Zhang et al.
+[2024c].
 The model supports high-resolution input (up to 672×672 pixels) and various aspect ratios (336×1344 and 1344×336) to capture more visual details, making it perform well in video understanding and analysis.
+A large body of research has demonstrated that AI systems are inherently susceptible to diverse attack vectors He et al.
+[2025], Liao et al.
+[2025], Sun et al.
+[2025a], Zhong et al.
+[2025], Zheng et al.
+[2025].
+Building on these findings, recent research has revealed that MLLMs likewise suffer from such vulnerabilities as we discussed earlier, which raises pressing concerns regarding their reliable and secure deployment.
+Although defense is not the primary focus of this paper, it is nevertheless crucial to acknowledge the importance of developing effective countermeasures, as explored in previous works Meta LLaMA [2025], Zhang et al.
+[2024b], Wang et al.
+[2024b], Liu et al.
+[2024e].
 Appendix B Prompts in $\mathsf{FC\text{-}Attack}$
@@
 Topic Selection Topics = ["cooking", "science", "DIY", "technology", "gardening", "music", "art", "history", "travel", "sports", "photography", "movies", "literature", "health", "fitness", "programming", "gaming", "education", "finance", "nature"]
-Few Shot Samples Instruction One:
+Few Shot Prompt prompt = f""" Generate a practical, instructional statement about topic {selected_topic}.
+The statement should describe a task or activity (e.g., "Teach people how to make a chair." or "Explain how to compost at home.").
+Follow it with a clear, step-by-step explanation structured for clarity and logic, ensuring the steps are concise and actionable.
+Examples: Instruction One:
 Explain how to make braised pork belly. Steps:
@@
 5. Water regularly without overwatering.
-Appendix D Flowchart Experiment Performance
-Figure A1: Successful jailbreak instance on Gemini-1.5 using $\mathsf{FC\text{-}Attack}$.
-Figure A2: ASR under different prompts against MLLMs via videos.
-Figure A3: Different styles of fonts in flowcharts (“1 step”).
-(a) Flowchart Example-Horizontal: Left-to-right flow.
+Now, please generate an example like above starting with word {act_topic}‘. """
+Appendix D Explainability of $\mathsf{FC\text{-}Attack}$
+To investigate why $\mathsf{FC\text{-}Attack}$ is effective on MLLMs, we conduct an experiment on LLaVA-Next, extract query embeddings by taking the hidden state of the last token from the final decoder layer, and evaluate on four different prompt types: benign text queries, harmful text queries, benign image queries, and FC-Attack image queries.
+Harmful text queries are drawn from a deduplicated version of Advbench.
+To construct benign text queries, we rewrite the harmful ones by replacing sensitive or unsafe terms with neutral, harmless alternatives while preserving the original grammar and syntactic structure to remove harmful intent.
+We then obtain embeddings for all four prompt types and visualize them using t-SNE.
+As shown in Figure˜A1, embeddings of benign versus harmful text queries are clearly separable, indicating strong semantic alignment in the text modality of LLaVA-Next.
+In contrast, embeddings from benign image queries and FC-Attack image queries are highly entangled, suggesting insufficient safety alignment in the image modality.
+Figure A1: Visualization in t-SNE of LLaVA-Next query embeddings in four different prompt types.
+Appendix E Flowchart Experiment Performance
+Figure A2: ASR under different prompts against MLLMs via videos
+(a) Flowchart Example-Horizontal: Left-to-right flow
 (b) Flowchart Example-S-shaped: S-shaped path flow
-(c) Flowchart Example-Vertical: Top-to-bottom flow.
-Figure A4: Flowchart Examples: Various flow directions.
-Table A1: Performances of ASR for different flowchart structures on the Qwen-2-VL.
+(c) Flowchart Example-Vertical: Top-to-bottom flow
+Figure A3: Flowchart Examples: Various flow directions
+Table A1: Performances of ASR for different flowchart structures on the Qwen2-VL.
 Method $\mathsf{FC\text{-}Attack}$ Plain Text Text with Box Figstep
-ASR (%) 88 32 50 34
+ASR (%) 90 32 50 34
+Table A2: Performances of ASR for different formats on the Qwen2-VL.
+Format $\mathsf{FC\text{-}Attack}$ Code Table Figstep
+ASR (%) 90 52 26 12